### v2

# Lorentzian Robin Universe of Gauss-Bonnet Gravity

Abstract
The gravitational path-integral of Gauss-Bonnet gravity is investigated and the transition from one spacelike boundary configuration to another is analyzed. Of particular interest is the case of Neumann and Robin boundary conditions which is known to lead to a stable Universe in Einstein-Hilbert gravity in four spacetime dimensions. After setting up the variational problem and computing the necessary boundary terms, the transition amplitude is computed exactly in the mini-superspace approximation. The $\hbar\to 0$ limit brings out the dominant pieces in the path-integral which is traced to an initial configuration corresponding to Hartle-Hawking no-boundary Universe. A deeper study involving Picard-Lefschetz methods not only allow us to find the integration contour along which the path-integral becomes convergent but also aids in understanding the crossover from Euclidean to Lorentzian signature. Saddle analysis further highlights the boundary configurations giving dominant contribution to the path-integral which is seen to be those corresponding to Hartle-Hawking no-boundary proposal and agrees with the exact computation. To ensure completeness, a comparison with the results from Wheeler-DeWitt equation is done.
KEYWORDS: Lorentzian Path-Integral, Boundary Conditions, Gauss-Bonnet Gravity, Picard-Lefschetz Methods, No-Boundary Universe, Wheeler-DeWitt Equation.
I Introduction
In quantum field theory, path integral is a useful tool for computing the transition amplitude from one initial configuration to a final one. It allows us to obtain physical information about the system, most importantly scattering cross sections/S-matrix. In flat/curved spacetime quantum field theory, the path integral computation involves the integration over the fields defined on the background of fixed flat/curved spacetime, obeying appropriate boundary conditions.
However, when gravity is present and/or background spacetime is dynamically evolving, as in the case of an expanding universe, defining a path-integral over fields is technically more involved. One has to carefully tackle issues like renormalizability, gauge-fixing, regularization, measure, boundary conditions and integration contour. Dealing with the last one comes with additional complications, which we aim to discuss here.
The pure gravitational path-integral (ignoring matter sector) with metric as the only degree of freedom can be schematically written as
$G[\rm Bd_{i},Bd_{f}]$ $=\int_{\mathcal{M}+\partial\mathcal{M}}\mathcal{D}g_{\mu\nu}e^{i\rm S[g_{\mu\nu}]/\hbar},$ (1)
$S[g_{\mu\nu}]$ $=\frac{1}{16\pi G_{\rm N}}\int d^{D}x\sqrt{-g}\left[-2\Lambda+R+\cdots\right]+\text{boundary terms}\,$ (2)
where $g_{\mu\nu}$ is the metric, $S[g_{\mu\nu}]$ is the corresponding gravity action, $R$ is the Ricci scalar, $G_{N}$ is Newton’s gravitational constant, $\Lambda$ is the cosmological constant, ()($\cdots$)( ⋯ ) includes the higher derivative curvature terms. The path integral is defined on manifold $\mathcal{M}$ with boundary $\partial\mathcal{M}$. $\rm Bd_{i}$ and $\rm Bd_{f}$ refer to field configurations at the initial and final boundaries, respectively.
The gravitational path-integral mentioned above in eq. (1) is hard to define suitably where meaningful computations can be done. The biggest trouble is the lack of renormalizability. Even if one keeps the issue of renormalizability aside, one still has to deal with the issue of boundary terms and integration contour, which is quite non-trivial as the standard flat spacetime Wick-rotation no longer works. To understand the importance and subtle nature of the last two bits (boundary terms and integration contour), one can reduce the gravitational phase-space by considering mini-superspace approximation. This reduced phase space gives us more freedom to play around with the gravitational path-integral, thereby allowing us to learn more about the boundary conditions and the integration contours.
In the mini-superspace approximation, the metric ansatz respecting spatially homogeneous and isotropic metric in $D$ spacetime dimensions is the FLRW metric. In polar coordinates $\{t_{p},r,\theta,\cdots\}$ it is given by
${\rm d}s^{2}=-N_{p}^{2}(t_{p}){\rm d}t_{p}^{2}+a^{2}(t_{p})\left[\frac{{\rm d}r^{2}}{1-kr^{2}}+r^{2}{\rm d}\Omega_{D-2}^{2}\right]\,.$ (3)
It has two unknown time-dependent functions: lapse $N_{p}(t_{p})$ and scale-factor $a(t_{p})$, $k=(0,\pm 1)$ is the scalar curvature, and ${\rm d}\Omega_{D-2}$ is the metric for the unit sphere in $D-2$ spatial dimensions. In this reduced space, although gravitational waves don’t exist, a reduced version of diffeomorphism invariance still shows in the time co-ordinate $t_{p}$ and the dynamical scale-factor $a(t_{p})$. This reduced setup is enough to explore the effects of boundary conditions and integration contour in the gravitational path-integral.
The path integral in this reduced gravitational phase space is
$G[{\rm Bd}_{f},{\rm Bd}_{i}]$ $=\int_{{\rm Bd}_{i}}^{{\rm Bd}_{f}}{\cal D}N_{p}{\cal D}\pi{\cal D}a(t_{p}){\cal D}p{\cal D}c{\cal D}\bar{c}{\cal D}\rho{\cal D}\bar{\rho}$ (4)
×$\quad\quad\times\exp\biggl{[}\frac{i}{\hbar}\int_{0}^{1}{\rm d}t_{p}\left(N^{\prime}_{p}\pi+a^{\prime}p+c^{\prime}\rho+\bar{c}^{\prime}\bar{\rho}-N_{p}H\right)\biggr{]}\,,$
where along with the integration over the scale-factor $a(t_{p})$, lapse $N_{p}(t_{p})$ the Fermionic ghost $c$ and the anti-ghost $\bar{c}$, we also have an integration over their respective conjugate momenta given by $p$, $\pi$, $\rho$, and $\bar{\rho}$ It should be specified that the $({}^{\prime})$ here denotes derivative with respect to $t_{p}$, and without any loss of generality the time co-ordinate is chosen to range from $0\leq t_{p}\leq 1$. ${\rm Bd}_{i}$ and ${\rm Bd}_{f}$ refers to the field configuration at initial ($t_{p}=0$) and final ($t_{p}=1$) boundaries respectively. The Hamiltonian constraint $H$ consists of two parts: gravitational part and ghost
$H=H_{\rm grav}[a,p]+H_{\rm gh}[N,\pi,c,\bar{c},\rho,\bar{\rho}]\,,$ (5)
where $H_{\rm grav}$ is the Hamiltonian for the gravitational action and the Batalin-Fradkin-Vilkovisky (BFV) Batalin:1977pb ghost Hamiltonian is denoted by $H_{\rm gh}$. BFV ghost generalizes the standard Fadeev-Popov ghost based on BRST symmetry. Unlike the gauge theories, in gravitational theories, the constraint algebra doesn’t close, thereby requiring BFV quantization. In minisuperspace, where we have only one constraint (Hamiltonian Constraint), the algebra trivially closes, but still, BFV quantization is preferable. Even though the phase space is quite reduced in the mini-superspace approximation, there is still time reparametrization invariance, which needs to be suitably gauge-fixed. A simple gauge choice is the ADM gauge also known as proper-time gauge $N^{\prime}_{p}=0$. More details on the BFV quantization process and BFV ghost can be found in the seminal works Teitelboim:1981ua ; Teitelboim:1983fk ; Halliwell:1988wc .
A large part of the above gravitational path-integral mentioned in eq. (4) can be performed exactly. This includes path-integral over the ghost-sector: $c,\bar{c}$ and $\rho,\bar{\rho}$; the conjugate momenta: $\pi$ and $p$. The path integral over the ghosts and the corresponding momenta leads to a lapse-independent constant term. Ignoring this constant term, the path-integral left behind is
$G[{\rm Bd}_{f},{\rm Bd}_{i}]=\int_{\bf C}{\rm d}N_{p}\int_{{\rm Bd}_{i}}^{{\rm Bd}_{f}}{\cal D}a(t_{p})\,\,e^{iS[a,N_{p}]/\hbar}\,.$ (6)
This path-integral can be interpreted as follows: $\int{\cal D}a(t_{p})\,\,e^{iS[a,N_{p}]/\hbar}$ gives the transition amplitude for the Universe to evolve from one boundary configuration to another in the proper time $N_{p}$, where ${\bf C}$ refers to the contour of Lapse integration. Depending on the choice of integration range ${\bf C}$ takes, eq. (6) leads to different objects. The $0<N_{p}<\infty$ choice leads to causal evolution from one boundary configuration ${\rm Bd}_{i}$ to another ${\rm Bd}_{f}$ as discussed in Teitelboim:1983fh , where $a_{0}<a_{1}$ will imply expanding Universe while $a_{0}>a_{1}$ will imply contracting Universe. Whereas the other choice, $-\infty<N_{p}<\infty$, takes the gauge invariance into account. For $-\infty<N_{p}<0$, $a_{0}<a_{1}$ will imply an expanding Universe while $a_{0}>a_{1}$ will imply contracting Universe where the subscript is with respect to $t_{p}$ coordinate. The $0<N_{p}<\infty$ leads to the (causal) propagator, while $-\infty<N_{p}<\infty$ leads to the wavefunction of the universe, which is a solution to the Wheeler-DeWitt equation. In this discussion, we focus on the latter case, with the lapse integration range $-\infty<N_{p}<\infty$. This choice allows us to account for all the dominant saddles in the path integral.
The gravitational action that we will focus on in this review article is given by following (see Narain:2021bff ; Narain:2022msz ; Ailiga:2023wzl for earlier works investigating the path integral of such gravitational theories)
$S=\frac{1}{16\pi G_{\rm N}}\int{\rm d}^{D}x\sqrt{-g}\biggl{[}-2\Lambda+R+\alpha\biggl{(}R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-4R_{\mu\nu}R^{\mu\nu}+R^{2}\biggr{)}\biggr{]}\,.$ (7)
This is the Gauss-Bonnet gravity action where $G_{\rm N}$ is the Newton’s gravitational constant, $\Lambda$ is the cosmological constant term, $\alpha$ is the Gauss-Bonnet (GB) coupling and $D$ is spacetime dimensionality. The mass dimensions of various couplings are: $[G_{N}]=M^{2-D}$, $[\Lambda]=M^{2}$ and $[\alpha]=M^{-2}$.
The gravitational action in eq. (7) falls in the class of lovelock gravity theories Lovelock:1971yv ; Lovelock:1972vz ; Lanczos:1938sf , and is a special type of higher-derivative gravity consisting of only two-time derivatives of metric-field. This has the privilege of keeping the dynamical evolution equation of the metric field second order in time. It should be pointed out that Gauss-Bonnet gravitational theories also arises in the low-energy effective action of the heterotic string theory with $\alpha>0$ Cheung:2016wjt ; Zwiebach:1985uq ; Gross:1986mw ; Metsaev:1987zx . Interestingly, this coupling also has received observational constraints Chakravarti:2022zeq coming from the study of the gravitational wave (GW) data of the event GW150914, which also offered the first observational confirmation of the area theorem Isi:2020tac .
In this short review article, we will focus on the path integral given in the eq. (6) for the gravitational action specified by eq. (7), and discuss the effects of the Neumann/Robin boundary conditions (NBC/RBC) on the transition amplitude (see York:1986lje ; Brown:1992bq ; Krishnan:2016mcj ; Witten:2018lgb ; Krishnan:2017bte for the role played by boundary terms). Imposing RBC at the initial boundary is like specifying a combination of scale factor and its conjugate momenta to be fixed. A particular class of the same is “linear” combination of scale factor and its conjugate momenta, characterized by a single parameter. These RBCs interpolate between Dirichlet BC and Neumann BC. Past studies involving specifying DBCs at the initial boundary showed unsuppressed growth of fluctuations around the saddles leading to instability in the no-boundary proposal of the Universe Feldbrugge:2017kzv ; DiTucci:2018fdg ; Lehners:2018eeo ; Feldbrugge:2017fcc ; Feldbrugge:2017mbc ; Matsui:2024bfn (see Lehners:2023yrj for review on the no-boundary proposal). These studies motivated us to investigate whether NBC or RBC can lead to stable behaviour of perturbations, also at the same time highlighting that Dirichlet BC is perhaps not the right boundary condition for gravity Krishnan:2016mcj ; Krishnan:2017bte .
Explorations with the Neumann boundary condition (NBC) have shown that gravitational perturbations are well-behaved in the no-boundary proposal of the Universe DiTucci:2019bui ; Narain:2021bff ; Lehners:2021jmv ; DiTucci:2020weq ; Narain:2022msz ; Mondal:2023cxx ; Ailiga:2024nkz . In the case of mini-superspace approximation with metric having ×$\mathbb{R}\times\mathbb{S}^{3}$ it is possible to do the computation of the path-integral exactly Narain:2022msz . These exact computations showed some interesting features where it was seen that our Universe undergoes signature changes from Euclidean to Lorentzian Narain:2022msz . Studies involving RBCs in Einstein-Hilbert gravity were done in DiTucci:2019dji ; Mondal:2022gyp ; DiTucci:2019bui under the WKB approximation using Picard-Lefschetz methods. In all these works it was noticed that perturbations were well-behaved. Exact computations imposing RBCs at the initial boundary have been done in Ailiga:2023wzl . Effects of Gauss-Bonnet addition have been thoroughly investigated in Narain:2021bff ; Narain:2022msz ; Ailiga:2023wzl .
This review also focuses on the impact of having Gauss-Bonnet addition in the gravitational action and the role it plays at the boundaries. This is crucial as the path-integral is sensitive to the boundary conditions. Retaining the consistency of the variational problem, the boundary condition is appropriately chosen. This still doesn’t guarantee that the solutions respecting these will be a stable configuration. But it is crucial to work those boundary conditions which lead to stable perturbations.
Such path-integrals need to be analyzed appropriately in a systematic manner in the arena of complex analysis. Earlier studies involving complex analysis methods to understand Euclidean gravitational path integral was done in Hawking:1981gb ; Hartle:1983ai . Boundary effects studies in the Euclidean quantum cosmology was done in tunnelling proposal made in Vilenkin:1982de ; Vilenkin:1983xq ; Vilenkin:1984wp and no-boundary proposal made in Hawking:1981gb ; Hartle:1983ai ; Hawking:1983hj . Later works investigating the contour choice in the path-integral involving the usage of complex analysis methods were also done in Halliwell:1988ik ; Halliwell:1989dy ; Halliwell:1990qr . More recent work Kontsevich:2021dmb on the ‘allowability’ criterion proposes a simple method to determine the allowable complex metric on which quantum field theories can be consistently defined. It is, however, yet to be understood and established whether this criterion is necessary or sufficient (see also recent works Witten:2021nzp and Lehners:2021mah ). Our approach avoids the standard process of Wick rotation used in the Euclidean approach and aims to directly tackle the Lorentzian gravitational path-integral using the Picard-Lefschetz (PL) method.
Picard-Lefschetz (PL) methodology generalizes the notion of standard Wick-rotation by suitably picking a contour along which oscillatory Lorentzian integrals can be performed in a reliable manner. It provides a framework that considers all the saddles of the path-integral and systematically picks only the ones which eventually lead to convergent path-integral. The original integration contour gets appropriately deformed into the integrals over the steepest descent contours passing through such saddles. Such contours termed Lefschetz thimbles together form the generalized Wick-rotated contour. In the Lorentzian quantum cosmology studies, these methods were utilized to understand the nature of the path integral Feldbrugge:2017kzv ; Feldbrugge:2017fcc ; Feldbrugge:2017mbc ; Vilenkin:2018dch ; Vilenkin:2018oja ; Rajeev:2021xit and the effects of the boundary conditions DiTucci:2019dji ; DiTucci:2019bui ; Narain:2021bff ; Ailiga:2024nkz ; Lehners:2021jmv ; Narain:2022msz ; Ailiga:2023wzl
The plan of this article is as follows: In section I, we present the introduction and discuss the motivation for studying the gravitational path integral. Section II discusses the mini-superspace approximation and outlines the surface terms required for the consistent variational problem. In section III, we compute the exact transition amplitude for Einstein-Gauss-Bonnet gravity with Neumann/Robin boundary conditions. We analyze the semiclassical behaviour along with the significance of Gauss-Bonnet in the section IV. In section IV.3, we discuss the significance of the Robin boundary condition. A brief overview of the Picard-Lefschetz technique is provided in the section V, followed by its application in the section VI to compute the saddle point results and verify their consistency with the results obtained from the exact analysis. Section VII outlines the connection between the path integral and canonical approaches. Finally, the article concludes with a summary and outlook in Section VIII.
II Mini-superspace action and Boundary Terms
The conformally flat FLRW metric mentioned in eq. (3) has vanishing Weyl-tensor $C_{\mu\nu\rho\sigma}=0$. The non-zero entries of the Reimann tensor for the $d$-dimensional FLRW is given by Deruelle:1989fj ; Tangherlini:1963bw ; Tangherlini:1986bw
$R_{0i0j}$ $=$ $-\left(\frac{a^{\prime\prime}}{a}-\frac{a^{\prime}N_{p}^{\prime}}{aN_{p}}\right)g_{ij}\,,$
$R_{ijkl}$ $=$ $\left(\frac{k}{a^{2}}+\frac{a^{\prime 2}}{N_{p}^{2}a^{2}}\right)\left(g_{ik}g_{jl}-g_{il}g_{jk}\right)\,,$ (8)
where $g_{ij}$ is the $D-1$-dimensional spatial part of the FLRW metric and $({}^{\prime})$ denotes derivative with respect to co-ordinate $t_{p}$. The non-zero entries of the Ricci-tensor for this metric are
$R_{00}$ $=$ $-(D-1)\left(\frac{a^{\prime\prime}}{a}-\frac{a^{\prime}N_{p}^{\prime}}{aN_{p}}\right)\,,$
$R_{ij}$ $=$ $\left[\frac{(D-2)(kN_{p}^{2}+a^{\prime 2})}{N_{p}^{2}a^{2}}+\frac{a^{\prime\prime}N_{p}-a^{\prime}N_{p}^{\prime}}{aN_{p}^{3}}\right]g_{ij}\,,$ (9)
while the Ricci-scalar for FLRW is given by
$R=2(D-1)\left[\frac{a^{\prime\prime}N_{p}-a^{\prime}N_{p}^{\prime}}{aN_{p}^{3}}+\frac{(D-2)(kN_{p}^{2}+a^{\prime 2})}{2N_{p}^{2}a^{2}}\right]\,.$ (10)
Interestingly, for conformal-flat (Weyl-flat) metrics, there is an identity relating the Riemann tensor with the Ricci-tensor and the Ricci scalar. This is given by,
$R_{\mu\nu\rho\sigma}=\frac{R_{\mu\rho}g_{\nu\sigma}-R_{\mu\sigma}g_{\nu\rho}+R_{\nu\sigma}g_{\mu\rho}-R_{\nu\rho}g_{\mu\sigma}}{D-2}-\frac{R(g_{\mu\rho}g_{\nu\sigma}-g_{\mu\sigma}g_{\nu\rho})}{(D-1)(D-2)}\,.$ (11)
By exploiting this identity one can express $R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}$ in terms of $R_{\mu\nu}R^{\mu\nu}$ and $R^{2}$ as follows
$R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}=\frac{4}{D-2}R_{\mu\nu}R^{\mu\nu}-\frac{2R^{2}}{(D-1)(D-2)}\,.$ (12)
This is a very useful relation when it is utilized in the GB-gravity action then, we get the following for case of Weyl-flat metrics
$\int{\rm d}^{D}x\sqrt{-g}$ $\left(R_{\mu\nu\rho\sigma}R^{\mu\nu\rho\sigma}-4R_{\mu\nu}R^{\mu\nu}+R^{2}\right)$ (13)
$=\frac{D-3}{D-2}\int{\rm d}^{D}x\sqrt{-g}\left(-4R_{\mu\nu}R^{\mu\nu}+\frac{DR^{2}}{D-1}\right)\,.$
The gravitational action stated in eq. (7) for the FLRW metric mentioned in eq. (3) is given by
$S=\frac{V_{D-1}}{16\pi G_{N}}\int{\rm d}t_{p}\biggl{[}\frac{a^{D-3}}{N_{p}^{2}}\biggl{\{}(D-1)(D-2)kN_{p}^{3}-2\Lambda a^{2}N_{p}^{3}-2(D-1)aa^{\prime}N_{p}^{\prime}$
$+(D-1)(D-2)a^{\prime 2}N_{p}+2(D-1)N_{p}aa^{\prime\prime}\biggr{\}}+(D-1)(D-2)(D-3)\alpha\biggl{\{}\frac{a^{D-5}(D-4)}{N_{p}^{3}}$
×$\times(kN_{p}^{2}+a^{\prime 2})^{2}+4a^{D-4}\frac{{\rm d}}{{\rm d}t_{p}}\left(\frac{ka^{\prime}}{N_{p}}+\frac{a^{\prime 3}}{3N_{p}^{3}}\right)\biggr{\}}\biggr{]}\,,$ (14)
where $V_{D-1}$ is the volume of $D-1$ dimensional sphere (for $\kappa=1$) and is given by  [^1] For $\kappa=(0,-1)$, the spatial slices are non-compact and $V_{D-1}$ wouldn’t be finite. In these cases, we have to either use an appropriate regularization scheme to regulate the volume or take non-trivial topologies.,
$V_{D-1}=\frac{\Gamma(1/2)}{\Gamma(D/2)}\left(\frac{\pi}{k}\right)^{(D-1)/2}\,.$ (15)
This is the action for scale-factor $a(t_{p})$ and lapse $N_{p}(t_{p})$ in $D$-dimensions. In $D=4$, something interesting happens in the terms proportional to $\alpha$. This term becomes total time-derivative. This is expected from Gauss-Bonnet gravity which is topological in $D=4$. The mini-superspace gravitational action then becomes the following in $D=4$
$S=\frac{V_{3}}{16\pi G_{N}}\int{\rm d}t_{p}\biggl{[}6kaN_{p}-2\Lambda a^{3}N_{p}-\frac{6a^{2}a^{\prime}N_{p}^{\prime}}{N_{p}}+\frac{6aa^{\prime 2}}{N_{p}}+\frac{6a^{\prime\prime}a^{2}}{N_{p}}+24\alpha\frac{{\rm d}}{{\rm d}t_{p}}\left(\frac{ka^{\prime}}{N_{p}}+\frac{a^{\prime 3}}{3N_{p}^{3}}\right)\biggr{]}\,.$ (16)
Lapse and scale-factor rescaling allows one to recast this action into a simpler form. This change of variables is given by
$N_{p}(t_{p}){\rm d}t_{p}=\frac{N(t)}{a(t)}{\rm d}t\,,\hskip 14.22636ptq(t)=a^{2}(t)\,.$ (17)
Then the original FLRW metric mentioned in eq. (3) changes into the following
${\rm d}s^{2}=-\frac{N^{2}}{q(t)}{\rm d}t^{2}+q(t)\left[\frac{{\rm d}r^{2}}{1-kr^{2}}+r^{2}{\rm d}\Omega_{D-2}^{2}\right]\,.$ (18)
Under this transformation, the gravitational action becomes the following
$S=\frac{V_{3}}{16\pi G_{N}}\int_{0}^{1}{\rm d}t\biggl{[}(6k-2\Lambda q)N+\frac{3\dot{q}^{2}}{2N}+3q\frac{{\rm d}}{{\rm d}t}\left(\frac{\dot{q}}{N}\right)+24\alpha\frac{{\rm d}}{{\rm d}t}\left(\frac{k\dot{q}}{2N}+\frac{\dot{q}^{3}}{24N^{3}}\right)\biggr{]}\,.$ (19)
In this the $(\dot{})$ represent time $t$-derivative. For some terms, one can do the integration by parts which allows us to simplify a bit more and combine some of the terms. This eventually leads to
$S$ $=\frac{V_{3}}{16\pi G_{N}}\int_{0}^{1}{\rm d}t\biggl{[}(6k-2\Lambda q)N-\frac{3\dot{q}^{2}}{2N}\biggr{]}+\frac{V_{3}}{16\pi G}\biggl{[}\frac{3q_{f}\dot{q}_{f}}{N}-\frac{3q_{i}\dot{q}_{i}}{N}$ (20)
$+24\alpha\left(\frac{k\dot{q_{f}}}{2N}+\frac{\dot{q_{f}}^{3}}{24N^{3}}-\frac{k\dot{q_{i}}}{2N}-\frac{\dot{q_{i}}^{3}}{24N^{3}}\right)\biggr{]}\,.$
Notice that there are two total derivative terms: one arises from the EH-part of gravitational action, while the other is the GB term itself which is topological in $4D$. We will keep $\Lambda$ to be arbitrary through the following sections till section IV, where we take it to be positive and do saddle point studies. Note also that in the following, we will make use of the convention $V_{3}=8\pi G_{N}$.
At this point, one can set up the variational problem by varying the action given in eq. (19) with respect to $q(t)$, which we do it by working in the ADM gauge $\dot{N}=0$ (setting $N(t)=N_{c}$, a constant). This can be implemented systematically by writing
$q(t)=\bar{q}(t)+\epsilon\delta q(t)$ (21)
where one assumes $\bar{q}(t)$ to satisfy the equation of motion, $\delta q(t)$ is the fluctuation around this. The parameter $\epsilon$ is inserted to keep track of terms in expansion, which becomes handy as one goes to higher orders in perturbative expansion.
$\delta S=\frac{\epsilon}{2}\int_{0}^{1}{\rm d}t\biggl{[}\left(-2\Lambda N_{c}+\frac{3\ddot{\bar{q}}}{N_{c}}\right)\delta q+\frac{3}{N_{c}}\frac{{\rm d}}{{\rm d}t}\left(\bar{q}\delta\dot{q}\right)+24\alpha\frac{{\rm d}}{{\rm d}t}\left\{\left(\frac{k}{2N_{c}}+\frac{\dot{q}^{2}}{8N_{c}^{3}}\right)\delta\dot{q}\right\}\biggr{]}\,.$ (22)
Note that in the above there are two total time-derivative terms. These eventually become boundary terms. For a consistent variational problem, these terms must either vanish for the given choice of boundary conditions or need to be appropriately cancelled by the addition of extra boundary terms. On the other hand, the collection of terms proportional to $\delta q$ gives the equation of motion for $q$,
$\ddot{\bar{q}}=\frac{2}{3}\Lambda N_{c}^{2}\,.$ (23)
The general solution to this equation is given by
$\bar{q}(t)=\frac{\Lambda N_{c}^{2}}{3}t^{2}+c_{1}t+c_{2}\,.$ (24)
The unknown parameters $c_{1,2}$ gets fixed by satisfying the appropriate boundary conditions. The collection of total-derivative terms is given by
$\delta S_{\rm bdy}=\frac{\epsilon}{2}\biggl{[}\frac{3}{N_{c}}\left(\bar{q}_{f}\delta\dot{q}_{f}-\bar{q}_{i}\delta\dot{q}_{i}\right)+24\alpha\left\{\left(\frac{k\delta\dot{q}_{f}}{2N_{c}}+\frac{\dot{\bar{q}}_{f}^{2}\delta\dot{q}_{f}}{8N_{c}^{3}}\right)-\left(\frac{k\delta\dot{q}_{i}}{2N_{c}}+\frac{\dot{\bar{q}}_{i}^{2}\delta\dot{q}_{i}}{8N_{c}^{3}}\right)\right\}\biggr{]}\,,$ (25)
where
[…]
The conjugate momenta corresponding to the field $q(t)$ can be computed from the action given in eq. (20)
$\pi=\frac{\partial{\cal L}}{\partial\dot{q}}=-\frac{3\dot{q}}{2N_{c}}\,,$ (27)
where we have used the ADM gauge. Utilizing the expression of the conjugate momentum, the boundary terms can be expressed in terms of $\pi$’s at the respective boundaries. This is given by
$\delta S_{\rm bdy}=-\epsilon\biggl{[}\left(\bar{q}_{f}\delta\pi_{f}-\bar{q}_{i}\delta\pi_{i}\right)+4\alpha\left\{\left(k\delta\pi_{f}+\frac{\bar{\pi}_{f}^{2}\delta\pi_{f}}{27}\right)-\left(k\delta\pi_{i}+\frac{\bar{\pi}_{i}^{2}\delta\pi_{i}}{27}\right)\right\}\biggr{]}\,.$ (28)
As mentioned before, for a consistent variational problem one has to get rid of these terms either by choosing boundary conditions where they vanish or by constructing suitable boundary action whose variation will cancel them exactly.
In the following, we will consider imposing Dirichlet boundary condition at $t=1$, but at $t=0$ we will consider two possibilities: Neumann or Robin boundary conditions, as they lead to suppressed perturbations Feldbrugge:2017fcc ; Feldbrugge:2017kzv .
II.1 Neumann Boundary condition (NBC) at $t=0$
In this section, we impose Neumann boundary condition Krishnan:2016mcj ; DiTucci:2019dji at $t=0$. This means the initial momentum and the final position are fixed,
[…]
Earlier studies have shown that this choice of boundary conditions lead to well-behaved perturbations which are suppressed DiTucci:2019bui ; Narain:2021bff ; DiTucci:2020weq ; Lehners:2021jmv . With the imposition of the above boundary condition, some of the terms in the eq. (28) will vanish. The residual terms need to be appropriately cancelled by addition of the surface terms whose variation will cancel them. The necessary surface terms are the following
.$\biggl{.}S_{\rm surface}\biggr{|}_{\rm NBC}=\frac{1}{2}\biggl{[}-\frac{3q_{f}\dot{q}_{f}}{N_{c}}-24\alpha\left(\frac{k\dot{q}_{f}}{2N_{c}}+\frac{\dot{q}_{f}^{3}}{24N_{c}^{3}}\right)\biggr{]}=q_{f}\pi_{f}+4\alpha\left(k\pi_{f}+\frac{\pi_{f}^{3}}{27}\right)\,.$ (30)
It should be mentioned here that the first term is the Gibbon-Hawking term for the mini-superspace while the second term a Chern-Simon like term at the final boundary. The constants $c_{1,2}$ appearing in the solution to the equation of motion eq. (24) can be now determined giving the solution $\bar{q}(t)$ to be
$\bar{q}(t)=\frac{\Lambda N_{c}^{2}}{3}(t^{2}-1)-\frac{2N_{c}\pi_{i}}{3}(t-1)+q_{f}\,,$ (31)
From this one can immediately note that $\bar{q}_{i}$ is related to $q_{f}$,
$\bar{q}_{i}=q_{f}+\frac{2N_{c}\pi_{i}}{3}-\frac{\Lambda N_{c}^{2}}{3}\,.$ (32)
This need not be true for off-shell configurations where $q_{i}$ could be anything. The surface terms when combined with the action in eq. (20) leads to the total action of the system. This is given by
$S_{\rm tot}[q,N_{c}]=\frac{1}{2}\int_{0}^{1}{\rm d}t\biggl{[}(6k-2\Lambda q)N_{c}-\frac{3\dot{q}^{2}}{2N_{c}}\biggr{]}+q_{i}\pi_{i}+4\alpha\left(k\pi_{i}+\frac{\pi_{i}^{3}}{27}\right)\,,$ (33)
Furthermore, if one plugs the solution to equation of motion in the above while also using the expression for $\bar{q}_{i}$ from eq. (32) one gets the total on-shell action
$S_{\rm tot}^{\rm on-shell}[\bar{q},N_{c}]=\frac{\Lambda^{2}}{9}N_{c}^{3}-\frac{\Lambda\pi_{i}}{3}N_{c}^{2}+\left(3k-\Lambda q_{f}+\frac{\pi_{i}^{2}}{3}\right)N_{c}+q_{f}\pi_{i}+4\alpha\left(k\pi_{i}+\frac{\pi_{i}^{3}}{27}\right)\,.$ (34)
This is also the action for the lapse $N_{c}$. It is important to note at this point that compared to the action obtained in the case of Dirichlet boundary condition, this doesn’t have singularity at $N_{c}=0$ Feldbrugge:2017kzv ; Feldbrugge:2017fcc ; Feldbrugge:2017mbc ; DiTucci:2019bui ; Narain:2021bff ; Lehners:2021jmv ; DiTucci:2020weq . The absence of singularity can be understood by realizing that in NBC path-integral as one doesn’t fix $q_{i}$ at $t=0$, transitions from all possible values of $q_{i}$ to a fixed $q_{f}$ are allowed, including the instantaneous transition occurring with $N_{c}=0$. This is only true for the Neumann BC and need not hold once the boundary conditions get generalized to Robin BC, in the sense that there will be a value of $N_{c}$ for which the action will be singular.
II.2 Robin Boundary condition (RBC) at $t=0$
Let’s now consider the case of Robin BC being imposed at the initial boundary with Dirichlet BC at the final boundary. This is a well posed boundary value problem where we fix the linear combination of $q_{i}$ and $\pi_{i}$ at the initial boundary, while keeping $q_{f}$ at the final boundary fixed,
$\pi_{i}+\beta q_{i}=P_{i}={\rm fixed}\hskip 8.53581pt\&\hskip 8.53581ptq_{f}={\rm fixed}\hskip 14.22636pt\Rightarrow\hskip 14.22636pt\delta P_{i}=0\hskip 8.53581pt\&\hskip 8.53581pt\delta q_{f}=0\,.$ (35)
When we have these boundary conditions, then the boundary terms mentioned in eq. (28) reduces to the following
.$\biggl{.}\delta S_{\rm bdy}\biggr{|}_{\rm RBC}=-\epsilon\biggl{[}\bar{q}_{f}\delta\pi_{f}+\beta\bar{q}_{i}\delta q_{i}+4\alpha\left\{k\delta\pi_{f}+\frac{\bar{\pi}_{f}^{2}\delta\pi_{f}}{9}+\beta\left(k\delta q_{i}+\frac{(\bar{P}_{i}-\beta\bar{q}_{i})^{2}\delta q_{i}}{9}\right)\right\}\biggr{]}\,.$ (36)
As before for consistent variational problem, we require that these terms either vanish or should be appropriately cancelled by the addition of suitable surface terms. It is seen that if one adds the following terms to the boundary
.$\biggl{.}S_{\rm surface}\biggr{|}_{\rm RBC}$ $=$ $q_{f}\pi_{f}+4\alpha\left(k\pi_{f}+\frac{\pi_{f}^{3}}{27}\right)+\frac{\beta}{2}(q_{f}^{2}+q_{i}^{2})$ (37)
$+4\alpha\beta\biggl{[}kq_{i}+\frac{q_{i}}{9}\left(P_{i}^{2}-\beta P_{i}q_{i}+\frac{\beta^{2}q_{i}^{2}}{3}\right)\biggr{]}\,,$
it leads to a consistent variational problem Ailiga:2023wzl . The first two terms here are the same as those obtained in the case of Neumann BC, while the last two terms appear additionally in the case of Robin BC. Moreover, the term in the last line arises when Gauss-Bonnet contribution is taken into account and was obtained for the first time in Ailiga:2023wzl . The additional terms are proportional to $\beta$ and vanish when $\beta\to 0$ (Neumann limit). The total action after combining these surface terms with the action in eq. (20) becomes
$S_{\rm tot}[q,N_{c}]=\frac{1}{2}\int_{0}^{1}{\rm d}t\biggl{[}(6k-2\Lambda q)N_{c}-\frac{3\dot{q}^{2}}{2N_{c}}\biggr{]}+q_{i}P_{i}+\frac{\beta}{2}(q_{f}^{2}-q_{i}^{2})+4\alpha\left(kP_{i}+\frac{P_{i}^{3}}{27}\right)\,.$ (38)
Once again in the limit $\beta\to 0$ this reduces to the total action for Neumann BC mentioned in eq. (33) as $P_{i}\to\pi_{i}$.
The constants appearing in the solution of the equation of motion $c_{1,2}$ can now be determined. This gives the solution
$\bar{q}(t)=\frac{\Lambda N_{c}^{2}}{3}t^{2}+\frac{P_{i}}{\beta}+\left(1+\frac{3}{2\beta N_{c}}\right)^{-1}\left(t+\frac{3}{2\beta N_{c}}\right)\left(q_{f}-\frac{P_{i}}{\beta}-\frac{\Lambda N_{c}^{2}}{3}\right)\,,$ (39)
Once again we note that on putting $t=0$ in the above we get a relation between $q_{i}$ and $q_{f}$
$\bar{q}_{i}=\frac{P_{i}}{\beta}+\left(\frac{3}{3+2\beta N_{c}}\right)\left(q_{f}-\frac{P_{i}}{\beta}-\frac{\Lambda N_{c}^{2}}{3}\right)\,.$ (40)
Off-shell $q_{i}$ can be anything. The on-shell total can now be worked out by making use of $\bar{q}(t)$ mentioned in eq. (39) into the action for RBC given in eq. (38). This is given by
$S_{\rm tot}^{\rm on-shell}[\bar{q},N_{c}]$ $=$ $\frac{1}{18(3+2N_{c}\beta)}\biggl{[}\beta\Lambda^{2}N_{c}^{4}+6\Lambda^{2}N_{c}^{3}+N_{c}^{2}\{108\beta k-18\Lambda(P_{i}+\beta q_{f})\}$ (41)
$+18N_{c}\left\{9k+P_{i}^{2}+q_{f}\left(\beta^{2}q_{f}-3\Lambda\right)\right\}+54P_{i}q_{f}\biggr{]}+4\alpha\left(kP_{i}+\frac{P_{i}^{3}}{27}\right).$
This is the lapse $N_{c}$ action in the RBC case. Note that in the limit $\beta\to 0$ this reduces to the lapse action for the NBC mentioned in eq. (34).
III Transition Amplitude
Having computed the total action for the lapse $N_{c}$, we proceed further to compute the transition amplitude from one three-dimensional spacelike hypersurface to another (for both NBC and RBC cases). The quantity of interest here (in the mini-superspace approximation) is as follows (see Halliwell:1988ik ; Feldbrugge:2017kzv )
$G[{\rm Bd}_{f},{\rm Bd}_{i}]=\int_{\bf C}{\rm d}N_{c}\int_{{\rm Bd}_{i}}^{{\rm Bd}_{f}}{\cal D}q(t)\,\,\exp\left(\frac{i}{\hbar}S_{\rm tot}[q,N_{c}]\right)\,,$ (42)
where as before ${\rm Bd}_{i}$ and ${\rm Bd}_{f}$ refers to initial and final boundary configurations respectively. The $S_{\rm tot}$ appearing above is the total action in the mini-superspace approximation. For the NBC case it is given in eq. (33) and for the RBC it is given in eq. (38). The NBC path-integral is not too hard to compute Narain:2022msz ; Ailiga:2023wzl . The RBC path-integral is a bit non-trivial, but can be computed exactly by making use of the procedure mentioned in section 2 of Ailiga:2023wzl .
The relevant expressions that will be needed in the following are the path-integral of a one-particle moving in linear potential whose action is given by
$S_{\rm tot}[q]=S[q]+S_{\rm bd}=\int_{0}^{1}{\rm d}t\left[\frac{m}{2}\dot{q}^{2}-\lambda q\right]+S_{\rm bd}\,,$ (43)
$m$ is a $t$-independent parameter and $\lambda$ is coupling parameter, $S_{\rm bd}$ is the surface term added to have a consistent variational problem. The path-integral we want to compute is
$\bar{G}[{\rm Bd_{f}},{\rm Bd_{i}}]=\int_{\rm Bd_{i}}^{\rm Bd_{f}}{\cal D}q(t)\,\,e^{iS_{\rm tot}[q]/\hbar}\,.$ (44)
For the initial Neumann and Robin boundary conditions they are given by the following expressions, respectively
$\bar{G}_{\rm NBC}(q_{f},t=1;p_{i},t=0)$ $=$ $\exp\biggl{[}\frac{i}{\hbar}\left\{p_{i}q_{f}-\frac{p_{i}^{2}}{2m}-\frac{\lambda\left(\lambda-3p_{i}+6mq_{f}\right)}{6m}\right\}\biggr{]}\,,$ (45)
$\bar{G}_{\rm RBC}(Q_{f},t=1;P_{i},t=0)$ $=$ $\sqrt{\frac{e^{i\beta Q_{f}^{2}}}{i\beta}}\int_{-\infty}^{\infty}\frac{{\rm d}\tilde{p}}{\sqrt{2\pi\hbar}}e^{i\frac{(P_{i}-\tilde{p})^{2}}{2\beta\hbar}}\bar{G}_{\rm NBC}(q_{f},t=1;\tilde{p},t=0)\,,$ (46)
where $p_{i}$ is the initial momentum, $q_{i}$ in initial position, $P_{i}=p_{i}+\beta\,q_{i}$ and $Q_{f}=q_{f}$ is the final position. The details of this derivation can be found in section 2 of Ailiga:2023wzl .
III.1 NBC at $t=0$
Let’s consider the mini-superspace path-integral mentioned in eq. (42) for the Neumann boundary condition. To compute this we make use of the expression given in eq. (45). In order to utilize this, we first compare the actions in eq. (43) with the NBC mini-superspace action given in eq. (33).
$m\to-\frac{3}{2N_{c}}\,,\hskip 28.45274ptV(q)=\lambda q\to\Lambda N_{c}q\hskip 8.53581pt\Rightarrow\hskip 8.53581pt\lambda\to\Lambda N_{c}\,,\hskip 28.45274ptp_{i}\to\pi_{i}\,.$ (47)
This gives the following expression for the Neumann boundary condition
$\int_{{\rm Bd}_{i}}^{{\rm Bd}_{f}}{\cal D}q(t)\exp\left(\frac{i}{\hbar}S_{\rm tot}[q,N_{c}]\right)=\exp\left(\frac{i}{\hbar}S_{\rm tot}^{\rm on-shell}[\bar{q},N_{c}]\right)$ (48)
$=$ $\exp\left[\frac{i}{\hbar}\left\{3kN_{c}+4\alpha\left(k\pi_{i}+\frac{\pi_{i}^{3}}{27}\right)\right\}\right]\bar{G}_{\rm NBC}[q_{f},t=1;\pi_{i},t=0]\,,$
where $S_{\rm tot}^{\rm on-shell}[\bar{q},N_{c}]$ is given in eq. (34) and $\bar{G}_{\rm NBC}[q_{f},t=1;\pi_{i},t=0]$ is given by
$\bar{G}_{\rm NBC}[q_{f},t=1;\pi_{i},t=0]=\exp\biggl{[}\frac{i}{\hbar}\biggl{\{}\frac{\Lambda^{2}N_{c}^{3}}{9}-\frac{\Lambda\pi_{i}N_{c}^{2}}{3}+\left(\frac{\pi_{i}^{2}}{3}-\Lambda q_{f}\right)N_{c}+\pi_{i}q_{f}\biggr{\}}\biggr{]}\,.$ (49)
It should be emphasized that the path-integral can also be performed using the zeta-function Narain:2022msz . However, the results differ from above by a numerical pre-factor. It is worth noting that the quantum fluctuation determinant for the Neumann boundary condition results in a constant, which is 1, that agrees with the Van Vleck-Pauli-Morette determinant (see doi:10.1142/7305 ). For a more detailed analysis, see Appendix A.
Note that $S_{\rm tot}^{\rm on-shell}[\bar{q},N_{c}]$ doesn’t posses a singularity at $N_{c}=0$ and the path-integral over $q(t)$ doesn’t give rise to additional singularities or branch-cuts. This is a feature of imposing Neumann BC at the initial boundary which doesn’t hold for a generic boundary condition. This allows to push the limit of $N_{c}$-integration all the way up to $\infty$- ∞ thereby giving
$G_{\rm NBC}[{\rm Bd}_{f},{\rm Bd}_{i}]=\int_{-\infty}^{\infty}{\rm d}N_{c}\,\,\exp\left(\frac{i}{\hbar}S_{\rm tot}^{\rm on-shell}[\bar{q},N_{c}]\right)\,,$ (50)
where $S^{\rm on-shell}_{\rm tot}[\bar{q},N_{c}]$ is given in eq. (34). By extending the range of integration all the way to $\infty$- ∞, one constructs the “ wave function ”, which is the solution of the WDW equation, as we later show in sec.VII. This is in contrast to the equation in 42, which computes the “propagator”. The integral can be done exactly by doing a change of variables achieved by shifting the lapse $N_{c}$ by a constant
[…]
Due to the change of variables the transition amplitude is given by
$G[{\rm Bd}_{f},{\rm Bd}_{i}]=$ $\frac{1}{2}\exp\left[\frac{i}{\hbar}\left(\frac{3}{\Lambda}+4\alpha\right)\left(k\pi_{i}+\frac{\pi_{i}^{3}}{27}\right)\right]$ (52)
×$\times\int_{-\infty}^{\infty}{\rm d}\bar{N}\,\,\exp\left[\frac{i}{\hbar}\left\{\frac{\Lambda^{2}}{9}\bar{N}^{3}+(3k-\Lambda q_{f})\bar{N}\right\}\right]=\Psi_{1}(\pi_{i})\Psi_{2}(q_{f})\,,$
where
$\Psi_{1}(\pi_{i})=\exp\left[\frac{i}{\hbar}\left(\frac{3}{\Lambda}+4\alpha\right)\left(k\pi_{i}+\frac{\pi_{i}^{3}}{27}\right)\right]\,,$ (53)
$\Psi_{2}(q_{f})=\int_{-\infty}^{\infty}{\rm d}\bar{N}\,\,\exp\left[\frac{i}{\hbar}\left\{\frac{\Lambda^{2}}{9}\bar{N}^{3}+(3k-\Lambda q_{f})\bar{N}\right\}\right]\,.$ (54)
It is seen that the transition amplitude with Neumann BC at the initial time and Dirichlet BC at the final time is a product of two parts: $\Psi_{1}(\pi_{i})$ and $\Psi_{2}(q_{f})$. The former is only dependent on the initial momentum, while the later is only a function of the final size of the Universe. This factorization has been observed in earlier studies of the Wheeler-DeWitt (WdW) equation in mini-superspace approximation of Einstein-Hilbert gravity Lehners:2021jmv (and also in Narain:2022msz ).
The function $\Psi_{2}(q_{f})$ can be computed as discussed in detail in the paper Narain:2022msz . It is proportional to Airy-function. This allow us to write a closed form expression for the transition amplitude in the NBC case as
$G_{\rm NBC}[{\rm Bd}_{f},{\rm Bd}_{i}]=$ $\sqrt{3}\left(\frac{3\hbar}{\Lambda^{2}}\right)^{\frac{1}{3}}\exp\left[\frac{i}{\hbar}\left(\frac{3}{\Lambda}+4\alpha\right)\left(k\pi_{i}+\frac{\pi_{i}^{3}}{27}\right)\right]$ (55)
×$\times Ai\left[\left(\frac{3}{\hbar^{2}\Lambda^{2}}\right)^{\frac{1}{3}}\left(3k-\Lambda q_{f}\right)\right]\,,$
III.2 RBC at $t=0$
We now focus on computing the expression in eq. (42) for the case of Robin BC at the initial boundary. For the Robin BC the $S_{\rm tot}[q,N_{c}]$ is given in eq. (38). Using the expression mentioned in eq. (46) one can relate the RBC path-integral with the NBC path-integral. This means we have
$\int_{{\rm Bd}_{i}}^{{\rm Bd}_{f}}$ ${\cal D}q(t)\,\,\exp\left(\frac{i}{\hbar}S_{\rm tot}[q,N_{c}]\right)=\exp\left[\frac{i}{\hbar}\left\{3kN_{c}+4\alpha\left(kP_{i}+\frac{P_{i}^{3}}{27}\right)\right\}\right]$ (56)
×$\times\left(\frac{2\pi\hbar}{i\beta}\right)^{1/2}\int_{-\infty}^{\infty}\frac{{\rm d}\tilde{p}}{2\pi\hbar}e^{i(P_{i}-\tilde{p})^{2}/2\hbar\beta}\bar{G}_{\rm NBC}[q_{f},t=1;\tilde{p},t=0]\,.$
This when plugged in the full path-integral involving also the lapse $N_{c}$-integration gives
$G_{\rm RBC}[{\rm Bd}_{f},{\rm Bd}_{i}]$ $=\left(\frac{2\pi\hbar}{i\beta}\right)^{1/2}\exp\left[\frac{4i\alpha}{\hbar}\left(kP_{i}+\frac{P_{i}^{3}}{27}\right)\right]\int_{-\infty}^{\infty}\frac{{\rm d}\tilde{p}}{2\pi\hbar}e^{i(P_{i}-\tilde{p})^{2}/2\hbar\beta}$ (57)
×$\times\int_{0^{+}}^{\infty}{\rm d}N_{c}\exp\left(\frac{3kN_{c}i}{\hbar}\right)\bar{G}_{\rm NBC}[q_{f},t=1;\tilde{p},t=0]\,.$
Note that the $N_{c}$-integrand is not singular at $N_{c}=0$ allowing us to push the limit of integration all the way up to $\infty$- ∞. This integral is however similar to the integral considered in the case of Neumann BC except for the absence of the Gauss-Bonnet contribution. This means one can write
$\int_{-\infty}^{\infty}{\rm d}N_{c}\exp\left(\frac{3kN_{c}i}{\hbar}\right)\bar{G}_{\rm NBC}[q_{f},t=1;\tilde{p},t=0]=\biggl{.}\Psi_{1}(\tilde{p})\biggr{|}_{\alpha=0}\times\Psi_{2}(q_{f})\,,$ (58)
where $\Psi_{1}$ and $\Psi_{2}$ are given in eq. (53) and (54) respectively. The $\Psi_{2}(q_{f})$ can be computed as in case of NBC and is proportional to Airy-function. The function $\Psi_{1}(\tilde{p})\biggr{|}_{\alpha=0}$ can also be computed by doing the following transformation
$\tilde{p}\to\bar{p}-\frac{3\Lambda}{2\beta}\,,$ (59)
which converts the original integral into an Airy-integral along with an exponential pre-factor. Putting all the pieces together we have in the RBC case
$G_{\rm RBC}[{\rm Bd}_{f},{\rm Bd}_{i}]=$ $\sqrt{\frac{6\pi\hbar}{i\beta}}\exp\left[\frac{4i\alpha}{\hbar}\left(kP_{i}+\frac{P_{i}^{3}}{27}\right)\right]\exp\biggl{[}\frac{i\bigl{(}-18k\beta^{2}+2P_{i}^{2}\beta^{2}+6P_{i}\beta\Lambda+3\Lambda^{2}\bigr{)}}{4\hbar\beta^{3}}\biggr{]}$ (60)
×$\times\left(\frac{9}{\Lambda\hbar}\right)^{\frac{1}{3}}Ai\left[\left(\frac{3}{\hbar^{2}\Lambda^{2}}\right)^{\frac{1}{3}}\left(3k-\Lambda q_{f}\right)\right]Ai\biggl{[}\left(\frac{3\Lambda}{\hbar^{2}}\right)^{\frac{1}{3}}\left(\frac{3k}{\Lambda}-\frac{P_{i}}{\beta}-\frac{3\Lambda}{4\beta^{2}}\right)\biggr{]}\,.$
It is worth noting that for the RBC case also, the path-integral once again factorizes into two parts: one depends only on initial parameters ($P_{i}$ and $\beta$), and another only on final $q_{f}$. This factorization was known previously to hold at the semiclassical level. However, at the exact expression, it was first shown in Ailiga:2023wzl .
For $\alpha>0$, a constraint motivated from Cheung:2016wjt ; Zwiebach:1985uq ; Gross:1986mw ; Metsaev:1987zx ; Chakravarti:2022zeq , it is noticed that the transition amplitude peaks around the initial parameter $P=-3i$, which is also the case, in absence of the Gauss-Bonnet for $\Lambda>0$. This result is illustrated in Figure 1, where we plot the term dependent on initial parameters, $G_{\rm RBC}(q_{f},P_{i})$, as a function of the Euclideanized variable $P_{i}=iy$. This favours the Hartle-Hawking no boundary geometry as the most dominant geometry in the path integral.
Figure 1: The plot of initial parameter dependent part of transition amplitude $G(q_{f},P_{i})$, given in Eq. (60). Here, we choose parameter values: $k=1$, $\Lambda=3$, $q_{f}=10$, $\hbar=0.0001$, $\alpha=1$ and $\beta=-1.5i$. The $G(q_{f},P_{i})$ axis has been scaled appropriately for ease of visualization.
In the subsequent section, we have shown it analytically. It is worth making the comparison with the NBC results in the limit $\beta\to 0$, in which limit the RBC reduces to NBC. To see this, we use the following identity
$\left(\frac{2\pi\hbar}{i\beta}\right)^{1/2}e^{i(P_{i}-\tilde{p})^{2}/2\hbar\beta}=\int_{-\infty}^{\infty}{\rm d}\xi\exp\left[-\frac{i\beta}{2\hbar}\xi^{2}+\frac{i}{\hbar}(P_{i}-\tilde{p})\xi\right]\,.$ (61)
Here, one immediately notices that in the limit $\beta\to 0$ one gets a $\delta$-function $\delta(P_{i}-\tilde{p})$. One can also obtain eq. 55 from the eq. 61 in the limit $\beta\rightarrow 0$, which involves asymptotic forms of the Airy-functions.
IV $\hbar\to 0$ limit
Having obtained the exact results for the transition amplitude, it is instructive to analyse the $\hbar\to 0$ limit. This is crucial as one needs to compare it with the results obtained from the saddle-point approximation using Picard-Lefschetz methods (to be discussed later on). This limit also brings out the configurations which contribute dominantly to the path-integral and those which are suppressed. In our context, this translates into the initial values for the $\pi_{i}$ for NBC, $P_{i}$ and $\beta$ for the RBC.
IV.1 Implications of having Gauss-Bonnet
Figure 2: Plot of Euclidean action $A_{E}$ given in eq. (64) vs $y$. The two saddle points are depicted by green and red squares. For $\alpha>0$: green one corresponds to $P_{i}=-3i\sqrt{k}$ (stable) while the red one corresponds to $P_{i}=3i\sqrt{k}$ (unstable). Ailiga:2023wzl
Let’s focus on the exponential prefactor arising from the Gauss-Bonnet sector, which reads,
$\exp\left[\frac{4i\alpha}{\hbar}\left(kP_{i}+\frac{P_{i}^{3}}{27}\right)\right]\equiv e^{iA(P_{i})/\hbar}\,,$ (62)
In the limit $\hbar\to 0$ the dominant contribution comes from geometries which extremize the exponent, which corresponds to the
$A^{\prime}(P_{i})=0\hskip 8.53581pt\Rightarrow\hskip 8.53581ptP_{i}=\pm 3i\sqrt{k}\,.$ (63)
In the limit $\beta\to 0$ (NBC case) this corresponds to $\pi_{i}=\pm 3i\sqrt{k}$! Two choices of signs correspond to two different choices for wick rotations. Imaginary momentum depicts the quantum nature of the initial condition. Let us consider taking $P_{i}=iy$, which is like considering a Euclideanised version action $A(P_{i})$. Then we have the Euclidean action $A_{E}(y)$ which is given by
$A_{E}(y)=4\alpha\left(ky-\frac{y^{3}}{27}\right)\,,\hskip 14.22636pte^{iA(P_{i})/\hbar}\hskip 8.53581pt\Rightarrow\hskip 8.53581pte^{-A_{E}(y)/\hbar}\,.$ (64)
On extremization of Euclidean action $A_{E}(y)$ we observe that for $\alpha>0$ the point $y=-3\sqrt{k}$ corresponds to a stable saddle point leading to a positive exponent (aka Hartle-Hawking ) while $y=3\sqrt{k}$ corresponds to unstable saddle point leading to exponential with a negative argument (Vilenkin Tunneling). Coincidently, as we have shown in IV.2 and was also mentioned in DiTucci:2019bui ; Narain:2022msz , $P_{i}=-3i\sqrt{k}$ also corresponds to initial condition giving regular and well-behaved geometries within a certain regime of $\beta$. Since a universe with a higher positive value of the exponent is more probable, the choice $P_{i}=-3i\sqrt{k}$ is favored, as it maximizes the Gauss-Bonnet exponent.
Although, technically allowing $\alpha<0$ would result in the point $P_{i}=3i\sqrt{k}$ becoming a dominant geometry in the path integral, which would lead to a Vilenkin Tunneling proposal. However, the studies in string theory, such as those in Cheung:2016wjt ; Zwiebach:1985uq ; Gross:1986mw ; Metsaev:1987zx , as well as observational constraints from gravitational wave data and the area theorem, discussed in Chakravarti:2022zeq ; Isi:2020tac , limit the value of $\alpha$ to positive values, restrict this possibility.
IV.2 Constraining $\beta$
Once $P_{i}$ gets fixed, the next thing that one has to do is to find the allowed possible values of parameter $\beta$ in the case of RBC transition amplitude. To find out the allowed range of $\beta$, one can analyse the nature of RBC transition amplitude in the $\hbar\to 0$ limit. This can be achieved by focusing only on the part of the wave function which depends only on initial configuration.
We start by plugging $P_{i}=-3i\sqrt{k}$ in eq. (60) and focusing first on the $\beta$-dependent Airy function. For $\beta$ negative imaginary ( as justified in section IV.3 ), the Airy function has a positive argument for $\Lambda>0$. This aids us in obtaining the $\hbar\to 0$ limit by utilizing the asymptotic structure of the Airy-function
$\left.Ai\biggl{[}\left(\frac{9\sqrt{\Lambda}}{\hbar}\right)^{2/3}\left(\sqrt{\frac{k}{\Lambda}}+\frac{i\sqrt{\Lambda}}{2\beta}\right)^{2}\biggr{]}\right|_{\hbar\to 0}\sim\exp\biggl{[}-\frac{6\sqrt{\Lambda}}{\hbar}\left|\sqrt{\frac{k}{\Lambda}}+\frac{i\sqrt{\Lambda}}{2\beta}\right|^{3}\biggr{]}=e^{-B_{1}(\beta)/\hbar}\,.$ (65)
We next focus on the $\beta$-dependent exponential factor in eq. (60) and study its nature for $P_{i}=-3i\sqrt{k}$ in $\hbar\to 0$ limit. For $P_{i}=-3i\sqrt{k}$ this exponential factor can be written as $e^{iB_{2}(\beta)/\hbar}$, where
$B_{2}(\beta)=\left(3\Lambda^{2}-18i\beta\Lambda\sqrt{k}-36k\beta^{2}\right)/(4\beta^{3})\,.$ (66)
Therefore, the overall exponential weighting for the RBC path integral in the semiclassical limit is expressed as:
$\exp\biggl{[}\frac{B(\beta)}{\hbar}\biggr{]}\equiv\exp\biggl{[}\frac{B_{1}(\beta)+B_{2}(\beta)}{\hbar}\biggr{]},$ (67)
From the structure, one could see that in limit $\hbar\rightarrow 0$ the dominant contribution comes from those configurations for which $B^{\prime}(\beta_{\rm dom})=0$. This means we have
$B^{\prime}(\beta_{\rm dom})=\frac{18i\Lambda}{\beta_{\rm dom}^{2}}\left(\sqrt{\frac{k}{\Lambda}}+\frac{i\sqrt{\Lambda}}{2\beta_{\rm dom}}\right)^{2}=0\,.$ (68)
This gives the dominant configuration when $\beta$ lies around $\beta_{\rm dom}=-i\Lambda/2\sqrt{k}$ (we don’t consider the $\beta_{\rm dom}\to\infty$ case). At this stage it is best to express
$\beta=-\frac{i\Lambda x}{2\sqrt{k}}=\beta_{\rm dom}x\hskip 14.22636pt{\rm where}\hskip 14.22636ptx\geq 0\,.$ (69)
This gives
$B_{1}(\beta)=6\frac{k^{3/2}}{\Lambda}\biggl{|}1-\frac{1}{x}\biggr{|}^{3}\,,\hskip 28.45274ptB_{2}(\beta)=-\frac{6ik^{3/2}\left(3x^{2}-3x+1\right)}{x^{3}\Lambda}\,.$ (70)
It immediately seen from here that there will be two cases in this: $0\leq x\leq 1$ and $x>1$. The dominant contributions comes from $x=1$ while $x=0$ refers to Neumann BC.
IV.2.1 $0\leq x\leq 1$
In this range of $x$ simplification occurs and various terms in the exponentials combine to give
$\exp\biggl{(}-\frac{B_{1}(\beta)}{\hbar}\biggr{)}\exp\left(\frac{iB_{2}(\beta)}{\hbar}\right)=\exp\biggl{[}\frac{6k^{3/2}}{\hbar\Lambda}\biggr{]}\,.$ (71)
When this is combined with the $\alpha$-dependent exponential pre-factor for $P_{i}=-3i\sqrt{k}$ (which is given by $\exp(8k^{3/2}\alpha/\hbar)$) one gets the following
$\exp\left(\frac{6k^{3/2}}{\hbar\bar{\Lambda}}\right)\hskip 8.53581pt\text{where,}\,\,\bar{\Lambda}=\Lambda/\left(1+4\alpha\Lambda/3\right)\,.$ (72)
This is the Hartle-Hawking state with positive weighting ($\Lambda>0$). The Gauss-Bonnet modification doesn’t change the overall structure of the exponent. This further supports the findings of the paper Jonas:2020pos , where it was noticed that Higher-derivative corrections don’t prohibit the no-boundary Universe as a solution. The analysis done in Jonas:2020pos was perturbative in nature. However, the topological nature of Gauss-Bonnet gravity helps us to go beyond the perturbative studies Ailiga:2023wzl . Although the contribution from Gauss-Bonnet appears as a multiplicative factor to the Einstein-Hilbert gravity, it is necessary to correctly reproduce the gravitation entropy of HDG theory Shu:2008yd . It is known that the H-H exponent is equal to entropy/2 in Einstein’s gravity. Our study also establishes this to be true for the higher-derivative gauss-bonnet gravity. Here, it is also seen that this exponential pre-factor is not only independent of the $\beta$-parameter but also hint that a lower value of $\bar{\Lambda}$ is more favourable.
IV.2.2 $x>1$
Things are a bit different in the case when $x>1$. In this situation we have
$\exp\biggl{(}-\frac{B_{1}(\beta)}{\hbar}\biggr{)}\exp\left(\frac{iB_{2}(\beta)}{\hbar}\right)=\exp\biggl{[}\frac{6k^{3/2}}{\hbar\Lambda}\left(-1+\frac{6}{x}-\frac{6}{x^{2}}+\frac{2}{x^{3}}\right)\biggr{]}\,.$ (73)
The total exponential factor also includes the factor coming from the Gauss-Bonnet sector $e^{8k^{3/2}\alpha/\hbar}$. It is easy to see that in the limit $x\to\infty$, we get inverse Hartle-Hawking. For any intermediate values $1<x<\infty$, the behaviour of exponential depends on the functional form of $x$. Suppose we call this function
$f(x)=-1+\frac{6}{x}-\frac{6}{x^{2}}+\frac{2}{x^{3}}\,.$ (74)
Then we notice that $f(1)=1$ and $f(\infty)=-1$. If we take derivative $f^{\prime}(x)=-6(x-1)^{2}/x^{4}<0$, then it is seen that it is always negative for all values of $x$ except $x=1$ and $x=\infty$. It is a monotonically decreasing function of $x$ in the range $1$ and $-1$, but at some point $x_{0}$ the function becomes zero. This $x_{0}=2+2^{1/3}+2^{2/3}$ is a crossover point. The function $f(x)>0$ for $1<x\leq x_{0}$, and $f(x)<0$ for $x>x_{0}$.
IV.3 Robin BC Interpretation
Before proceeding further, let us comment on a few things regarding why the Robin boundary condition at $t=0$ might be useful in the context of H-H no boundary proposal, which is a direct consequence of the above choice of $\beta$ (purely imaginary). In this case, one can interpret the boundary terms as being coherent state i.e.,
$\exp\left[i\left(-\frac{\beta q_{i}^{2}}{2}+P_{i}q_{i}\right)\right]\xrightarrow[]{\text{for,\,\,}\beta=-i|\beta|}\exp\left[-\frac{|\beta|q_{i}^{2}}{2}+iP_{i}q_{i}\right]\,.$ (75)
This is a coherent state with an imaginary momentum $P_{i}$ with the peak at $q_{i}=0$. One must be a little bit careful before making such an interpretation. Substituting naively $P_{i}=-ip$ with $p$ being real in eq. 75, one ends up with a real state with zero (mean) momentum with a peak at $q_{i}=p/|\beta|$. However, we should view the state in eq. 75 as a complex generalization of the usual coherent state in the (analytically continued) phase space ($P_{i},q_{i}$). With this in mind, eq.75 describes a state with shared uncertainty between the scale factor and the momentum of the universe at $t=0$. Taking this as an initial state DiTucci:2019bui ; DiTucci:2019dji , one can express the wave function as a gravitational path-integral in the following way
[…]
where $S_{DD}$ is action with the Dirichlet BC at the two endpoints. In the limit $\lvert\beta\rvert\to\infty$, $e^{-i\beta q_{i}^{2}/2}\to\delta(q_{i})$. This is a strict imposition of the Dirichlet BC $q_{i}=0$. The other limit $\lvert\beta\rvert\to 0$ gives Neumann BC for which the state becomes a simple plane wave. Finite values of $\beta$ can be seen as a form of ’regulator’, embedding a regularized version of $\delta$-function in the path-integral. In a sense, RBC is a regularised DBC. This boundary condition is also compatible with the more fundamental quantum uncertainty principle in the sense knowing the initial size ($q_{i}$) arbitrarily accurate would render the initial momentum ($P_{i}$) completely undetermined or vice versa. For example, the Neumann boundary condition at $t=0$ will allow the universe to start with an arbitrarily large size, which seems problematic to the no boundary philosophy. Moreover, quantum uncertainty doesn’t allow to determine the initial size (and/, or initial momentum) with arbitrary accuracy DiTucci:2019dji . In a way, studying Robin BC is the most appropriate due to its compatibility with the uncertainty principle and the regularized way of introducing Dirichlet BC overcoming technical challenges associated with dealing delta function in path integrals.
V Picard-Lefschetz Methodology: mini review
The mini-superspace path-integral mentioned in eq. (42) involves an integration over the real lapse $N_{c}$. Both in the NBC and RBC case one has to workout this integral either by directly computing it (as in NBC) or converting it into an integral over $\tilde{p}$ (as in RBC ). These integrals can also be worked out in saddle-point approximation and using Picard-Lefschetz methods. These methods helps us in analysing the behavior of the integrand in the complex plane Witten:2010cx ; Witten:2010zr ; Basar:2013eka ; Tanizaki:2014xba . This is achieved by determining the set of saddle points in the complex plane and finding the associated steepest ascent/descent paths. Then the Picard-Lefschetz methods systematically tell us the deformed integration contour along which the integrand is well-behaved and integral is convergent. To understand this in more detail, we consider a generic oscillatory integral
[…]
where the exponent is a function of $z$ (real). Here $\hbar$ is a real parameter, and the action ${\cal S}(z)$ is a real-valued function. Generically, such integrals are not absolutely convergent along the given contour. It may be either divergent or conditionally convergent. However, if it is conditionally convergent the Picard-Lefschetz (PL) methodology gives a systematic procedure for determining the deformed contour of integration along which the integral becomes absolutely convergent. In PL theory, one analytically continues both the integration parameter $z$ and the function ${\cal S}(z)$ into the complex plane. ${\cal S}(z)$ is then interpreted as an holomorphic function of $z$. This implies that ${\cal S}$ satisfies the Cauchy-Riemann conditions
$\frac{\delta{S}}{\delta\bar{z}}=0\Rightarrow\begin{cases}\frac{\delta{\rm Re}{S}}{\delta x_{1}}&=\frac{\delta{\rm Im}{S}}{\delta x_{2}}\,,\\ \frac{\delta{\rm Re}{S}}{\delta x_{2}}&=-\frac{\delta{\rm Im}{S}}{\delta x_{1}}\,.\end{cases}$ (78)
V.1 Flow equations
If one writes the complex exponent as ${\cal I}=i{\cal S}/\hbar=h+iH$, and expresses $z(\lambda)=x_{1}(\lambda)+ix_{2}(\lambda)$, then the downward flow can be written as
$\frac{{\rm d}x_{i}}{{\rm d}\lambda}=-g_{ij}\frac{\partial h}{\partial x_{j}}\,,$ (79)
where the metric $g_{ij}$ appearing above is a Riemannian metric defined on the complex plane, $\lambda$ is the flow parameter, and the ()(-)( - ) sign represents the downward flow. The solution to these equations gives the steepest descent trajectories associated with the corresponding saddle point (also known as critical points). These are also known as thimbles (denoted by ${\cal J}_{\sigma}$, $\sigma$ subscript symbolizes saddles). If, in the above set of equations, we have a (+)(+)( + ) sign instead of ()(-)( - ), then we have the equations for the steepest ascent flow lines (denoted by ${\cal K}_{\sigma}$). The flow parameter ($\lambda$) takes the values in the range of ($\infty$- ∞,+$\infty$+ ∞). For a downward flow emanating from the saddle, $\lambda$ takes $\infty$- ∞ at the saddles and increases as we move along the flow line away from the saddles. Along the steepest ascent line, $\lambda$ increases as we move towards the saddle point and takes +$\infty$+ ∞ at the saddles. The subscript $\sigma$ refers to the saddle point to which these sets of trajectories are attached. From this definition, it is quickly seen that the real part $h$ (also called Morse function) decreases monotonically along the steepest descent lines while increasing monotonically along the steepest ascent ones. To confirm this one can consider
$\frac{{\rm d}h}{{\rm d}\lambda}=g_{ij}\frac{{\rm d}x^{i}}{{\rm d}\lambda}\frac{\partial h}{\partial x_{j}}=-\left(\frac{{\rm d}x_{i}}{{\rm d}\lambda}\frac{{\rm d}x^{i}}{{\rm d}\lambda}\right)\leq 0\,.$ (80)
This holds true for any generic Riemannian metric (matrices which have a positive inner product). However, for simplicity and understanding the nature of these equations, it is useful to adopt a metric (Kahler metric). In our case, we’ll opt for the flat Cartesian metric, $ds^{2}=\lvert dx\rvert^{2}$. With complex coordinates $(z,\bar{z}=(x+iy,x-iy))$, the metric components are $g_{z,z}=g_{\bar{z},\bar{z}}=0$ and $g_{z,\bar{z}}=g_{\bar{z},z}=1/2$. This simplifies flow equations as:
$\frac{{\rm d}z}{{\rm d}\lambda}=\pm\frac{\partial\bar{\cal I}}{\partial\bar{z}}\,,\hskip 14.22636pt\frac{{\rm d}\bar{z}}{{\rm d}\lambda}=\pm\frac{\partial{\cal I}}{\partial z}\,.$ (81)
From this set of equations, one notices that the imaginary part of ${\rm Im}{\cal I}=H$ is constant along the flow lines.
[…]
This feature can be exploited to find the flow trajectories, which otherwise maybe hard to find out as one has to deal with the differential equations. However, in order to find the nature of trajectory (whether steepest ascent/descent), one just needs to analyse the behaviour of ${\rm d}h(\lambda)/{\rm d}\lambda$. Note that for real ${\cal S}(z)$, the morse function $h$ will vanish on the real line. This means that $h$ along the steepest ascent trajectories which intersect the real line will be more negative when traced backwards. Hence the saddle points whose associated steepest ascent line intersects the real line (original integration contour) will have negative $h$. Interestingly, the integration done along the thimbles of this saddle will be convergent by definition. These saddles are called relevant. In a generic situation, when $S(z)$ is not necessarily real to start with, only those saddles will be called relevant from which the associated steepest ascent lines intersect the original integration contour. These saddles contribute to the gravitational path integral and hence are called relevant.
V.2 Choice of contour
Once the saddle points and associated flow lines have been computed, one needs to work out a way to deform the original integration contour and make the integral absolutely convergent (for more detail see Witten:2010zr ; Basar:2013eka ; Feldbrugge:2017kzv ).
In the complex plane, the behaviour of $h$ and $H$ will aid us in determining the correct deformed contour of integration. If one denotes a particular saddle as $N_{\sigma}$ then the region is called ‘allowed’ if for each point $N$ in the region we have $h(N)<h(N_{\sigma})$. This ‘allowed’ region is denoted by $J_{\sigma}$. A region is called ‘forbidden’ if, for each point $N$ in the region, we have $h(N)>h(N_{\sigma})$. The ‘forbidden’ region is denoted by $K_{\sigma}$. Intuitively, this can be thought of as a landscape ridded with valleys (allowed region), hills (forbidden region) and plains (intermediate region).
It is also important to note that $h(\lambda)\to-\infty$ along the steepest descent lines ($\mathcal{J}_{\sigma}$) when $\lambda\to+\infty$ (i.e., moving away from the saddles), while $h(\lambda)\to\infty$ along the steepest ascent ($\mathcal{K}_{\sigma}$) lines when $\lambda\to-\infty$ (i.e., moving away from the saddles). This implies that the integrand is dying out as we move away from the saddles along the steepest descent thimbles, and the integral along the whole thimble is dominated by the saddle point contribution only. This is what we need to validate the saddle point approximation. However, along the steepest ascent lines, the integrand increases exponentially as we move away from the saddle points, and this invalidates the saddle point approximation. These lines ($\mathcal{J}_{\sigma}$ and $\mathcal{K}_{\sigma}$) intersect only at the saddle point and where they satisfy222 We have already removed the degeneracies related to the stokes ray which connects two saddle points. This happens when the steepest descent curve from a saddle point meets the steepest ascent curve from another saddle point as a result of symmetry. This can be removed by adding a small perturbation (off-shell fluctuation) to the exponent.
${\rm Int}\left({\cal J}_{\sigma},{\cal K}_{\sigma^{\prime}}\right)=\delta_{\sigma\sigma^{\prime}}\,,$ (83)
where “$\rm Int$” counts the intersection between two curves. Now let us deform the original integration contour ($\mathbb{D}$) to the deformed contour ($\mathcal{C}$) and write $\mathcal{C}=\sum_{\sigma}n_{\sigma}\mathcal{J}_{\sigma}$, for some integers $n_{\sigma}$ which will take value 00 or ±$\pm 1$ depending on the orientation. This gives $n_{\sigma}={\rm Int}({\cal C},{\cal K}_{\sigma})={\rm Int}(\mathbb{D},{\cal K}_{\sigma})$. Being topological $n_{\sigma}$ doesn’t change if one deforms the contours. It also implies that the necessary and sufficient condition for a saddle to be relevant if the steepest ascent emanating from the saddle intersects the original integration contour for which only $n_{\sigma}$ is non-zero. The corresponding thimble ${\cal J}_{\sigma}$ becomes the deformed integration contour. It is also crucial that the deformed integration contour shouldn’t lie in the forbidden region near the endpoints of the original integration, which we always keep fixed during the deformation. This ensures that we can safely slide the original integration contour into the deformed integration contour by cauchy’s theorem.
It is often the case that more than one saddle is relevant, and hence there will be multiple thimbles that one would need to integrate over. This will imply that the original integral has been converted into integrals over all the relevant thimbles
$I=\int_{\cal C}{\rm d}ze^{i\mathcal{S}(z)/\hbar}=\sum_{\sigma}n_{\sigma}\int_{{\cal J}_{\sigma}}{\rm d}ze^{i\mathcal{S}(z)/\hbar}\,.$ (84)
This is the feature of the Picard-Lefschetz methodology. The integration done along each of the thimble is absolutely convergent as can be seen from
$\biggl{|}\int_{{\cal J}_{\sigma}}{\rm d}ze^{i\mathcal{S}(z)/\hbar}\biggr{|}\leq\int_{{\cal J}_{\sigma}}\lvert{\rm d}z\rvert\lvert e^{i\mathcal{S}(z)/\hbar}\rvert=\int_{{\cal J}_{\sigma}}\lvert{\rm d}z\rvert e^{h}(z)<\infty\,.$ (85)
Denoting the ‘distance’ along the contour path as $l=\int\lvert{\rm d}z\rvert$. Then for the above integral to converge it is necessary that $e^{h}\sim 1/l$ as $l\to\infty$. Consequently, the original integration can be analytically deformed into a sum of absolutely convergent integrals along various thimbles passing through the relevant saddle points. By expanding in $\hbar$, the leading-order expression is as follows:
$I=\int_{\cal C}{\rm d}ze^{i\mathcal{S}(z)/\hbar}=\sum_{\sigma}n_{\sigma}e^{iH(z_{\sigma})}\int_{{\cal J}_{\sigma}}{\rm d}ze^{h}\approx\sum_{\sigma}n_{\sigma}e^{i\mathcal{S}(z_{\sigma})/\hbar}\left[{\cal A}_{\sigma}+{\cal O}(\hbar)\right]\,.$ (86)
Here, ${\cal A}_{\sigma}$ represents the contribution (fluctuation) obtained after performing a Gaussian integration around the saddle point $z_{\sigma}$.
VI Saddle point approximation
Having detoured a bit to learn about the Picard-Lehschetz, we return back to analyzing the mini-superspace path-integral. To proceed further and gain deeper understanding of the results obtained so far it is best to do saddle-point studies and compare it with the exact results and those obtained under the $\hbar\to 0$ limit. The $\hbar\to 0$ study has shown us that the dominant contribution in the path-integral comes from the $P_{i}=\pm 3i\sqrt{k}$, and only $P_{i}=-3i\sqrt{k}$ correspond to stable configuration. This is also the value of $P_{i}$, which leads to stable behaviour of perturbations in the Hartle-Hawking no-boundary proposal of the Universe DiTucci:2019bui ; Narain:2021bff ; Lehners:2021jmv ; DiTucci:2020weq ; DiTucci:2019dji ; Narain:2022msz . It is important to highlight that in a sense the Gauss-Bonnet favours the Hartle-Hawking no-boundary Universe.
We will now do the saddle point analysis of the of the mini-superspace path-integral for the RBC case (we will skip the NBC analysis which is presented in Narain:2022msz ). However, we will not apply the Picard-Lefschetz methods directly on the $N_{c}$-integral, instead as the RBC path-integral can be written as an NBC path-integral using eq. (57) and (58), one can then study the integral over the $\tilde{p}$ and $N_{c}$ using PL separately.
In the integral over $N_{c}$ mentioned in eq. (58) if one makes a shift as in eq. (51) then it separates in two parts: one dependent on $\tilde{p}$ and other an integral over $\bar{N}$ (not dependent on $\tilde{p}$). The $\bar{N}$-integral depends on $q_{f}$ is written as $\Psi_{2}(q_{f})$, and its saddle analysis has been done in Narain:2022msz (won’t be included here). It was noticed that for all values of $q_{f}$ there are two saddle points whose nature changes for $q_{f}<3k/\Lambda$, $q_{f}=3k/\Lambda$ and $q_{f}>3k/\Lambda$. In the first case both saddles lie on imaginary axis however only one of them is relevant. In the case when $q_{f}>3k/\Lambda$, both saddle are complex in nature and both are relevant. The case of $q_{f}=3k/\Lambda$ is a degenerate case. For $q_{f}<3k/\Lambda$, Universe is Euclidean as can be seen from the exponential growth of the transition amplitude referring to imaginary time, while in the $q_{f}>3k/\Lambda$ the transition amplitude is oscillatory in nature indicating the emergence of real time and a Lorentzian Universe.
We then move to analysing the behaviour of the $\tilde{p}$-integral. This integral is given by
$\int_{-\infty}^{\infty}\frac{{\rm d}\tilde{p}}{2\pi\hbar}e^{i(P_{i}-\tilde{p})^{2}/2\hbar\beta}\biggl{.}\Psi_{1}(\tilde{p})\biggr{|}_{\alpha=0}=\int_{-\infty}^{\infty}\frac{{\rm d}\tilde{p}}{2\pi\hbar}e^{iB(\tilde{p})/\hbar}=\int_{-\infty}^{\infty}\frac{{\rm d}\tilde{p}}{2\pi\hbar}e^{\{h(\tilde{p})+iH(\tilde{p})\}/\hbar}\,,$ (87)
where
$B(\tilde{p})=\frac{(P_{i}-\tilde{p})^{2}}{2\beta}+\frac{3}{\Lambda}\left(k\tilde{p}+\frac{\tilde{p}^{3}}{27}\right)\,.$ (88)
and $h(\tilde{p})$ is the corresponding Morse-function while the $H(\tilde{p})$ corresponds to real part of the $B$-function. Once again one can recast this into an Airy-integral by a shift of variable as discussed previously. Here we study this using Picard-Lefschetz methods (see Feldbrugge:2017kzv ; Narain:2021bff ; Witten:2010cx ; Witten:2010zr ; Basar:2013eka ; Tanizaki:2014xba for review on Picard-Lefschtez and analytic continuation).
The saddle-points of $\tilde{p}$ can be obtained by by setting ${\rm d}B(\tilde{p})/{\rm d}\tilde{p}=0$. These are given by
$\tilde{p}_{1}=-3i\sqrt{k}\,,\hskip 14.22636pt\tilde{p}_{2}=\frac{3i\sqrt{k}(x-2)}{x}\,.$ (89)
This shows that while saddle point $\tilde{p}_{1}$ remains fixed at the same position for all $x$, the saddle point $\tilde{p}_{2}$ moves as $x$ changes. It becomes zero for $x=2$. At the saddle point $\tilde{p}_{1}$, we have a vanishing initial on-shell geometry, i.e., $\bar{q}_{i}=0$, while at the other saddle point $\tilde{p}_{2}$, we have a non-vanishing initial geometry $\bar{q}_{i}=12k(x-1)/\Lambda x^{2}$.
One can expand the function $B(\tilde{p})$ around the saddle point $\tilde{p}_{\sigma}$ where $\sigma=\{1,2\}$. This gives
$B(\tilde{p})=B(\tilde{p}_{\sigma})+\biggl{.}\frac{{\rm d}B(\tilde{p})}{{\rm d}\tilde{p}}\biggr{|}_{\tilde{p}_{\sigma}}\delta\tilde{p}+\frac{1}{2}\biggl{.}\frac{{\rm d}^{2}B(\tilde{p})}{{\rm d}\tilde{p}^{2}}\biggr{|}_{\tilde{p}_{\sigma}}(\delta\tilde{p})^{2}+\frac{1}{6}\biggl{.}\frac{{\rm d}^{3}B(\tilde{p})}{{\rm d}\tilde{p}^{3}}\biggr{|}_{\tilde{p}_{\sigma}}(\delta\tilde{p})^{3}\,,$ (90)
where $\delta\tilde{p}=\tilde{p}-\tilde{p}_{\sigma}$. The second variation at the saddle-point can be written as $B^{\prime\prime}(\tilde{p}_{\sigma})=r_{\sigma}e^{i\rho_{\sigma}}$, where $r_{\sigma}$ and $\rho_{\sigma}$ depends on boundary conditions. Near the saddle point, the change in $iH$ will go like
$\delta(iH)\propto i\left(B^{\prime\prime}(\tilde{p}_{\sigma})\right)\left(\delta\tilde{p}\right)^{2}\sim v_{\sigma}^{2}e^{i\left(\pi/2+2\theta_{\sigma}+\rho_{\sigma}\right)}\,,$ (91)
where we write $\delta\tilde{p}=v_{\sigma}e^{i\theta_{\sigma}}$ and $\theta_{\sigma}$ is the direction of flow lines at $\tilde{p}_{\sigma}$ and given by
$\theta_{\sigma}=\frac{(2k-1)\pi}{4}-\frac{\rho_{\sigma}}{2}\,,$ (92)
where $k\in\mathbb{Z}$.
For the steepest descent and ascent flow lines, their corresponding $\theta_{\sigma}^{\rm des/aes}$ is such that the phase for $\delta H$ correspond to $e^{i\left(\pi/2+2\theta_{\sigma}+\rho_{\sigma}\right)}=\mp 1$. This implies
[…]
These angles can be computed numerically in our case.
Under the saddle point approximation, the contour integral given in eq. (87) can be computed using Picard-Lefschetz methods (see Narain:2021bff for details of PL methodology). This gives
$\int_{-\infty}^{\infty}\frac{{\rm d}\tilde{p}}{2\pi\hbar}e^{iB(\tilde{p})/\hbar}$ $=\sum_{\sigma}\frac{n_{\sigma}}{2\pi\hbar}e^{iB(\tilde{p}_{\sigma})/\hbar}\int_{{\cal J}_{\sigma}}{\rm d}\delta\tilde{p}\exp\left[i\frac{B^{\prime\prime}(\tilde{p}_{\sigma})}{2\hbar}(\delta\tilde{p})^{2}\right]\,,$ (94)
$=\sum_{\sigma}n_{\sigma}\sqrt{\frac{1}{2\pi\hbar\lvert B^{\prime\prime}(\tilde{p}_{\sigma})\rvert}}e^{i\theta_{\sigma}}e^{iB(\tilde{p}_{\sigma})/\hbar}\,.$
where ${\cal J}_{\sigma}$ refers to steepest descent line/thimble and $n_{\sigma}$ is the intersection number which will take values ($0,\pm 1$) accounting for the orientation of contour over each thimble. In the following we will be use this expression to compute the contour integral for various values of $x$.
VI.1 $10<x<$ 1
For $10<x<$ 1, both the saddle points $\tilde{p}_{1}$ and $\tilde{p}_{2}$ lie on the negative imaginary axis with $\lvert\tilde{p}_{2}\rvert<\lvert\tilde{p}_{1}\rvert$.
Figure 3: This figure depicts the Picard-Lefschetz analysis of the contour integral given in eq. (87). The red lines correspond to the steepest descent lines (thimbles ${\cal J}_{\sigma}$), while the thin black lines are the steepest ascent lines and are denoted by ${\cal K}_{\sigma}$. Here, we choose parameter values: $k=1$, $\Lambda=3$, and $x=1/2$. For this, the saddle point $\tilde{p}_{1}$ corresponds to the blue square, and it is relevant, while saddle point $\tilde{p}_{2}$ corresponds to the blue circle. The thick black line shows the original integration contour (,+)($\infty$,$\infty$)( - ∞ , + ∞ ). The light-green region is the allowed region with $h<h(\tilde{p}_{\sigma})$ for all values of $\sigma$. The light-orange region (forbidden region) has $h>h(\tilde{p}_{\sigma})$ for all $\sigma$. The intermediate region is depicted in yellow Ailiga:2023wzl .
In figure 3, we plot the various flow-line, saddle points, and forbidden/allowed regions. From the graph, we notice that only the steepest ascent line from $\tilde{p}_{1}$ intersects the original integration contour (real line), thereby making it relevant. This contour picks the correct Airy-function to be $Ai(.)$. At this saddle point, the initial geometry turns out to be vanishing, thereby satisfying both the “compactness” and “regularity” criterion of Hartle-Hawking’s no-boundary proposal. The Picard-Lefschetz theory then gives the following in the saddle point approximation as
$\int_{-\infty}^{\infty}\frac{{\rm d}\tilde{p}}{2\pi\hbar}e^{i(P_{i}-\tilde{p})^{2}/2\hbar\beta}\biggl{.}\Psi_{1}(\tilde{p})\biggr{|}_{\alpha=0}\approx\sqrt{\frac{\Lambda xk^{-1/2}}{4\hbar\pi(1-x)}}\,\,e^{\frac{6k^{3/2}}{\hbar\Lambda}}\,.$ (95)
Hence, we get the same exponential factor as in the Hartle-Hawking no-boundary Universe. Gauss-Bonnet being topological in 4D won’t alter this as long as perturbations are small.
VI.2 $x>1$
For $x>1$, only saddle point $\tilde{p}_{1}$ remains fixed on the negative imaginary axis for all values of $x$, while the saddle point $\tilde{p}_{2}$ moves from negative imaginary axis to positive imaginary axis.
Refer to caption           Refer to caption
Figure 4: Here we choose parameter values: $k=1$, $\Lambda=3$, and $x=4$ (left) and $x=1$ (right) respectively. For the left figure, the saddle point $\tilde{p}_{2}$ corresponds to blue- circle, while saddle point $\tilde{p}_{1}$ corresponds to blue-square. Only the saddle point $\tilde{p}_{2}$ (blue square) is relevant. The right figure corresponds to the degenerate case when two saddle points merge, and they are both relevant.
We see that only the steepest ascent curves emanating from $\tilde{p}_{2}$ intersect the original integration contour, implying that it is the only relevant saddle point. At this saddle, the initial geometry of the universe is non-vanishing, given by $\bar{q}_{i}=12k(x-1)/{\Lambda x^{2}}$. The descent thimble passing through this saddle will be the deformed contour of integration, which again picks $Ai(.)$ in eq. (87). The Picard-Lefschetz analysis then gives the following in the saddle point approximation as
$\int_{-\infty}^{\infty}\frac{{\rm d}\tilde{p}}{2\pi\hbar}e^{i(P_{i}-\tilde{p})^{2}/2\hbar\beta}\biggl{.}\Psi_{1}(\tilde{p})\biggr{|}_{\alpha=0}\approx\sqrt{\frac{\Lambda xk^{-1/2}}{4\pi\hbar(x-1)}}\,\,e^{\frac{6k^{3/2}}{\hbar\Lambda}\left(\frac{2}{x^{3}}-\frac{6}{x^{2}}+\frac{6}{x}-1\right)}\,.$ (96)
VI.3 $x=1$
This is the degenerate case where both the saddle points merge : $\tilde{p}_{1}=\tilde{p}_{2}=-3i\sqrt{k}$, with vanishing initial geometry at the relevant saddle. In this case the Morse-function becomes $h(\tilde{p}_{1})=h(\tilde{p}_{2})=6k^{3/2}/\Lambda$. Saddle-point approximation breaks down as the second derivative vanishes, $B^{\prime\prime}(\tilde{p}_{\sigma})=0$. One needs to go third order in the series expansion in eq. (90). This situation is depicted in figure 4. In this case, the integration gives the following
$\int_{-\infty}^{\infty}\frac{{\rm d}\tilde{p}}{2\pi\hbar}e^{i(P_{i}-\tilde{p})^{2}/2\hbar\beta}\biggl{.}\Psi_{1}(\tilde{p})\biggr{|}_{\alpha=0}\approx\frac{\sqrt{3}}{2\pi\hbar}e^{iB(\tilde{p}_{\sigma})/\hbar}\int_{0}^{\infty}{\rm d}v_{\sigma}e^{-v_{\sigma}^{3}/9\Lambda\hbar}=\frac{(\Lambda/3\hbar^{2})^{1/3}}{\Gamma(2/3)}e^{\frac{6k^{3/2}}{\hbar\Lambda}}\,.$ (97)
VII Connection to WDW equation
So far, we have been focusing on the path integral quantization. Employing the path integral approach is both technically and conceptually advantageous, especially when the initial condition is specified on the geometry, such as in this case, on the metric variable, rather than directly on the wave-function (as in the case of the DeWitt condition). However, one can ask how these results are related to the solution of the Wheeler-DeWitt equation in the canonical approach.
For the mini-superspace model eq. (18), the Lagrangian is given by:
$L=-\frac{3}{4N}\dot{q}^{2}+3N-N\Lambda q\,,$ (98)
where we have set $k=1$ and done integration by parts in eq. (19) to obtain the above Lagrangian. The canonical momentum $p$ conjugate to $q$ is:
$p=\frac{\partial L}{\partial q}=-\frac{3}{2N}\dot{q}.$ (99)
Consequently, the Hamiltonian takes the following form
${H}_{\rm{EH}}=\dot{q}p-L=N\left[-\frac{p^{2}}{3}+(\Lambda q-3k)\right]=N\hat{H}_{EH}\,.$ (100)
The Gauss-Bonnet Hamiltonian in 4 spacetime dimensions is zero Liu:2008zf i.e., ${H}_{GB}=0$. This is because the Gauss-Bonnet term is topological in four dimensions and should not affect the bulk dynamics.
The Hamiltonian describing the Einstein-Gauss-Bonnet model is given by
$\hat{H}_{EGB}=\hat{H}_{\rm{EH}}+\hat{H}_{\rm{GB}}=\hat{H}_{\rm{EH}}\,.$ (101)
The Wheeler-DeWitt equation corresponds to the operator version of classical constraint $H_{EGB}=0$:
$\hat{H}_{EGB}\Psi=0\,,$ (102)
where $p$, $q$ appearing in eq. (100) have been promoted to operators with the commutation relation $[\hat{q},\hat{p}]=i\hbar$. Eq. 102 implies that physical states are annihilated by the Hamiltonian.
In the following, we will employ this equation at $t=0$, where we impose the robin boundary condition $p_{i}+\beta q_{i}=P_{i},q_{i}=Q_{i}$. Consequently, the Hamiltonian constraint takes the form
$\left[-\frac{1}{3}\left(P_{i}-\beta Q_{i}\right)^{2}+\left(\Lambda Q_{i}-3\right)\right]=0.$ (103)
In order to derive the Wheeler-DeWitt (WDW) equation, we promote the variables to operators. It will be convenient to transition into momentum space by substituting $\hat{Q}_{i}=i\hbar\frac{\partial}{\partial P_{i}}$ and $\hat{P}_{i}=P_{i}$. Then, the WDW equation reads
$\frac{1}{3}\left(P_{i}^{2}\Psi_{0}+i\hbar\beta P_{i}\frac{\partial\Psi_{0}}{\partial P_{i}}+i\hbar\beta\frac{\partial}{\partial P_{i}}(P_{i}\Psi_{0})-\beta^{2}\hbar^{2}\frac{\partial^{2}\Psi_{0}}{\partial P_{i}^{2}}\right)+\left(i\Lambda\hbar\frac{\partial\Psi_{0}}{\partial P_{i}}+3\Psi_{0}\right)=0$ (104)
where $\Psi_{0}=\Psi_{0}(P_{i})$. Note that in eq. 104, we flip the sign $\frac{\partial}{\partial P_{i}}\rightarrow-\frac{\partial}{\partial P_{i}}$ because we are considering this at the initial boundary. It is a second-order differential equation. The most general solution is given by [^3] The Airy functions $Ai$ and $Bi$ are defined as two linearly independent solutions. Rotating the arguments by $32\pi/$ 3 results different linear combinations of Airy functions, given by: $Ai(ze^{\pm i\frac{2\pi}{3}}=\frac{1}{2}e^{\pm i\frac{\pi}{3}}\left[Ai(z)\mp iBi(z)\right]$:
$\begin{split}\Psi_{0}(P_{i})=\exp\left[i\left(\frac{P_{i}^{2}}{2\hbar\beta}+\frac{3P_{i}\Lambda}{2\hbar\beta^{2}}\right)\right]\times\left[c_{1}Ai\biggl{[}\left(\frac{-3\Lambda}{\hbar^{2}}\right)^{\frac{1}{3}}\left(-\frac{3}{\Lambda}+\frac{P_{i}}{\beta}+\frac{3\Lambda}{4\beta^{2}}\right)\biggr{]}\right.\\ +c_{2}\left.Bi\biggl{[}\left(\frac{-3\Lambda}{\hbar^{2}}\right)^{\frac{1}{3}}\left(-\frac{3}{\Lambda}+\frac{P_{i}}{\beta}+\frac{3\Lambda}{4\beta^{2}}\right)\biggr{]}\right]\end{split}$ (105)
where, $c_{1}$ and $c_{2}$ are arbitrary functions of $\beta$ and $\Lambda$. To match with the $t=0$ part of the path-integral expression in eq. 60, we choose the following,
$\begin{split}c_{1}=-ic_{2}=\frac{1}{2}\sqrt{\frac{3\pi\hbar}{2i\beta}}e^{-i\pi/3}\exp\left[i\left(-\frac{9}{2\hbar\beta}+\frac{3\Lambda^{2}}{4\hbar\beta^{3}}\right)\right]\end{split}$ (106)
This fixes the constants $c_{1}$ and $c_{2}$ uniquely. At this stage, it’s worth noting a basic difference between $\beta=0$ (Neumann condition) and $\beta\neq 0$ cases. For $\beta=0$, the Wheeler-DeWitt (WDW) equation is a first-order differential equation and hence has only one solution. In this scenario, the WDW equation provides the correct wave-function without any reference to the path-integral. The reason behind this is that computing the $t=0$ part from the path-integral doesn’t require any specific choice of contour. The lesson learned is that the specific choice of contour directly relates to the arbitrary constants generated in the WDW equation. At $t=1$, it’s convenient to work in position space by substituting $p_{f}=-i\hbar\frac{\partial}{\partial q_{f}}$. Then the WDW equation reads
$\frac{\hbar^{2}}{3}\frac{\partial^{2}\Psi_{1}}{\partial q_{f}^{2}}+(\Lambda q_{f}-3)\Psi_{1}=0,$ (107)
where $\Psi_{1}=\Psi_{1}(q_{f})$. The general solution is given by
$\Psi_{1}(q_{f})=b_{1}Ai\left[\left(\frac{-\sqrt{3}}{\Lambda\hbar}\right)^{\frac{2}{3}}\left(3-q_{f}\Lambda\right)\right]+b_{2}Bi\left[\left(\frac{-\sqrt{3}}{\Lambda\hbar}\right)^{\frac{2}{3}}\left(3-q_{f}\Lambda\right)\right]$ (108)
As before, to match with the $t=1$ part of the path integral expression in eq. 60, we choose the following,
$b_{1}=-ib_{2}=\frac{1}{2}e^{-i\pi/3}$ (109)
This fixes the constants $b_{1}$, $b_{2}$ uniquely at $t=1$. Combining $t=0$ and $t=1$ parts together, we get the total wave-function $\Psi(P_{i},q_{f})=\Psi_{0}(P_{i})\Psi_{1}(q_{f})$. Hence, path integral indeed generates the solutions of the WDW equation for all values of initial momentum $P_{i}$ and $\beta$ (alternatively, for a given $P_{i}$, it is a solution for all $\beta$ and vice-versa.). The exponential pre-factor in eq. 60 coming from the topological Gauss-Bonnet ($\alpha$) part cannot be captured via the WDW equation. However, if we demand that the full wave function, given in eq. (60) obtained from the path-integral approach, keeping the contribution of $\alpha$ to satisfy the WDW equation, which is certainly a more stringent criterion, we notice that only for $P_{i}=-3i\sqrt{k}$ the WDW equation is satisfied at the semi-classical level. To see this, we will assume $P_{i}=-iy,\beta=-i\frac{\Lambda}{2}x$, which is the relevant range of the parameters to make the total wave-function real. Using these relabellings, the WDW equation reads in the semi-classical limit:
$0\displaystyle 4\alpha\left(k+\frac{P_{i}^{2}}{9}\right)\left[\frac{\alpha\Lambda x}{3k}\left(k+\frac{P_{i}^{2}}{9}\right)+\left(1-\frac{2iP_{i}}{3x\sqrt{k}}+\frac{1}{x^{2}}\right)^{1/2}\right]=$ 0 . (110)
Now, assume $P_{i}\neq\pm 3i\sqrt{k}$, then eq. 110 will be satisfied if the term inside the square bracket vanishes. Since the $\alpha$ dependent term also contains $\Lambda$, upon vanishing, this will make $\alpha$ and $\Lambda$ dependent at arbitrarily chosen $(x,y)$ point. This certainly doesn’t make sense as $\alpha$ and $\Lambda$ are two independent parameters of our theory. Alternatively, if we take $\alpha$ and $\Lambda$ independently, then the eq. 110 equation is satisfied for a given $y$, only for specific $x$ values, which is not what we wanted. On the other hand, for $P_{i}=\pm 3i\sqrt{k}$, the eq. 110 satisfied identically for all values of $\alpha$, $\Lambda$ and $x$, which makes sense and is along the line of our expectation. From this argument, we conclude allowing Gauss-bonnet picks “only” specific values of $P_{i}$ which implement no boundary. Hence, eliminating all solutions except those no-boundary solutions. The last statement supports that no-boundary solutions give the dominant contribution to the gravitational path integral at the semi-classical level.
Now, we would like to analyse the no-boundary criterion (vanishing $q$) from the canonical perspective, which reads
$\hat{Q}_{i}\Psi_{0}(P_{i})=0\,\,\rightarrow\,\,\frac{\partial\Psi_{0}}{\partial P_{i}}=0.$ (111)
Substituting $\Psi_{0}(P_{i})$ and then taking the derivative, we get after simple algebra in the semi-classical limit the following constraint
$\left(y-\frac{3}{x}\right)+\left(9-\frac{6y}{x}+\frac{9}{x^{2}}\right)^{\frac{1}{2}}=0\,,$ (112)
where we substituted $P_{i}=-iy$ and $\beta=-i\frac{\Lambda}{2}x$. From eq. 112, we get $y=3$ and $0\leq x\leq 1$ for the Hartle-Hawking no-boundary Universe agreeing with the previous findings. Similarly to the case of the WDW equation, demanding the entire wave function given in eq. (60), including the Gauss-Bonnet contribution, to satisfy the above no-boundary criterion:
$\hat{Q}_{i}G_{\rm RBC}=0\,\,\rightarrow\,\,\frac{\partial G_{\rm RBC}}{\partial P_{i}}=0\,.$ (113)
We obtain the semi-classical constraint as:
$\alpha\Lambda x\left(1-\frac{y^{2}}{9}\right)+\left(y-\frac{3}{x}\right)+\left(9-\frac{6y}{x}+\frac{9}{x^{2}}\right)^{\frac{1}{2}}=0,$ (114)
Remarkably, again the solution $y=3$ in the interval $0\leq x\leq 1$ identically satisfies the above constraint for all values of $\alpha$ and $\Lambda$, thereby realizing the Hartle-Hawking no-boundary proposal. This demonstrates that the conclusions derived from the canonical perspective are aligned with those obtained from the path integral approach, even in the presence of the Gauss-Bonnet term.
We note that this discussion may well justify the choice of measure used in the path integral in previous sections.
VIII Discussion and outlook
In this article, we investigate the gravitational path integral of Einstein-Gauss-Bonnet gravity in four spacetime dimensions. Our study focuses on the Lorentzian path integral within the mini-superspace approximation. We extensively examine the effects of imposing the Robin boundary condition (fixing the linear combination of scale factor and its conjugate momentum) at the initial time, and the Dirichlet boundary condition (fixing the scale factor) at the final time and investigate the role played by the Gauss-Bonnet sector of gravity
We carefully construct the surface terms needed for the Einstein-Gauss-Bonnet gravity with Robin boundary condition at an initial time and Dirichlet at the final, and supplement to the action to ensure a consistent variational problem. The path-integral is then studied with this total action.
Utilizing the relations connecting the Robin boundary condition (RBC) path integral as an integral transform of the Neumann BC path-integral, and drawing upon the results of past studies Narain:2022msz , where the exact expression for the NBC path-integral has been derived, we can obtain the exact expression for the RBC path integral. This result is a product of two Airy functions, where one is a function of the initial parameters $P_{i}$ $\beta$, and the other is a function of the final parameter $q_{f}$, with an exponential pre-factor that also depends on the initial parameters alone. An important observation regarding the path-integral is that the term dependent on the Gauss-Bonnet coupling ($\alpha$) appears only in the exponential pre-factor. When put together, all this gives the exact transition amplitude in the RBC case, which is given in eq. (60). In the limit $\beta\to 0$ this correctly reproduces the exact NBC transition amplitude, which is given in eq. (55). This exact transition amplitude for the RBC including the Gauss-Bonnet effects, was first derived in Ailiga:2023wzl .
We then examined the nature of transition amplitude in the limit $\hbar\to 0$ (also known as the semi-classical limit). This allowed us to identify and focus on configurations that contribute dominantly to the path-integral. Primarily, the Gauss-Bonnet contribution, which appears as an exponential pre-factor only shows that in the $\hbar\to 0$ limit, two configurations give dominant contribution: $P_{i}=-3i\sqrt{k}$ and $P_{i}=3i\sqrt{k}$. The former corresponds to a stable configuration, while the later is unstable. Interestingly, $P_{i}=-3i\sqrt{k}$ coincides with the configuration for the Hartle-Hawking no-boundary Universe, where perturbations are well-behaved. In a sense, the Gauss-Bonnet contribution naturally picks and favours the Hartle-Hawking no-boundary condition.
Later, we investigated the $\hbar\to 0$ limit of other terms in the RBC transition amplitude for the case $P_{i}=-3i\sqrt{k}$ (Hartle-Hawking no-boundary), a choice deeply motivated by the stability of perturbations associated with this configuration. This analysis revealed that $\beta=\beta_{\rm dom}=-i\Lambda/2\sqrt{k}$ yields the dominant contribution. Upon rescaling $\beta=\beta_{\rm dom}x$, we observed that for $x\leq 1$, we obtain the same Hartle-Hawking exponential pre-factor $e^{6k^{3/2}/\hbar\bar{\Lambda}}$, where $\bar{\Lambda}$ is given in eq. (72). For $x>1$, the argument of the exponential pre-factor monotonically decreases from positive weighting to negative weighting, asymptotically approaching $e^{-\frac{6k^{3/2}}{\hbar\Lambda}\left(1-\frac{4\alpha\Lambda}{3}\right)}$. This analysis shows that the allowed region where we correctly reproduce the Hartle-Hawking exponential prefactor is $0\leq x\leq 1$.
We then proceeded to study the saddle point analysis of the $\tilde{p}$ integral 87. Making use of the Picard-Lefschetz technique, we evaluated the oscillatory integral and obtained the amplitude expression in the saddle point approximation, which indeed matches with the semiclassical results of exact transition amplitude. The saddle analysis of the $\tilde{p}$-integration reveals two saddle points in the complex $\tilde{p}$ plane, both lying on the imaginary axis. The saddle point $\tilde{p}_{1}$ remains fixed, independent of $x$, with a vanishing initial geometry, while the saddle point $\tilde{p}_{2}$ varies with respect to $x$, representing a non-vanishing initial geometry, and tends towards infinity (on the negative imaginary axis) as $x$ approaches 0 (Neumann limit). For all values of $x$, only one of the two saddle points is relevant. For $10<x<$ 1, the saddle point $\tilde{p}_{1}$ is relevant. In this regime, we obtain the correct Hartle-Hawking exponential pre-factor $e^{6k^{3/2}/\hbar\bar{\Lambda}}$. For $x>1$, the saddle $\tilde{p}_{2}$ becomes relevant. At $x=1$, there is a degenerate situation where $\tilde{p}_{1}=\tilde{p}_{2}$ (both relevant), resulting in $e^{6k^{3/2}/\hbar\bar{\Lambda}}$ (Hartle-Hawking) with additional numerical pre-factors. Thus, the range of $x$ for which only the saddle point $\tilde{p}_{1}$ is relevant and produces the exact exponential pre-factor $e^{6k^{3/2}/\hbar\bar{\Lambda}}$ of the Hartle-Hawking no-boundary Universe is $10<x<$ 1.
In the subsequent part of the article, we explored the relationship between the results obtained through the path integral approach and those derived from the canonical approach. Although, the topological Gauss-Bonnet ($\alpha$) part cannot be captured via the Wheeler-DeWitt (WDW) equation. However, when demanding that the full wave function, as given in eq. (60), satisfies the WDW equation, it led to the conclusion that this condition is only satisfied for the values $P_{i}=\pm 3i\sqrt{k}$ for all values of $\alpha$. Also, when demanding the entire wave function to satisfy the no-boundary criterion from a canonical perspective, we find that $P_{i}=-3i\sqrt{k}$ in the limit $0\leq x\leq 1$ is the configuration for which it is satisfied for all values of $\alpha$. Remarkably, these findings align with the results obtained from the path integral approach, where $P_{i}=\pm 3i\sqrt{k}$ were identified as the dominant contributions in the semiclassical limit for all values of $\alpha$. Moreover, $P_{i}=-3i\sqrt{k}$ and $0\leq x\leq 1$ is the configuration for which the Hartle-Hawking no-boundary proposal is realized.
Our investigations show the important non-trivial role played by the Gauss-Bonnet sector of the gravitational action in favouring the initial configurations which lead to the Hartle-Hawking no-boundary Universe. Gauss-Bonnet term arises in the low Energy-effective action of the heterotic string theory Cheung:2016wjt ; Zwiebach:1985uq ; Gross:1986mw ; Metsaev:1987zx with $\alpha>0$. Although it is topological in nature in four spacetime dimensions and is expected to not play any role in the dynamical evolution, but our analysis clearly shows its contribution to the path integral. When Gauss-Bonnet term is not ignored in path-integral studies, then for $\alpha>0$ (which is also the same sign appearing in low energy effective action of the heterotic string theory) it naturally picks and favors initial configurations which correspond to Hartle-Hawking no-boundary Universe. Moreover, Gauss-Bonnet modifications won’t alter the stability analysis done for Einstein-Hilbert gravity DiTucci:2019dji . This is because Gauss-Bonnet in four spacetime dimension is topological. Perturbative analysis like the one done in DiTucci:2019dji (see also Ailiga:2024nkz ) is for small perturbations which are expected to not change the background topology, and hence will be unaffected by the Gauss-Bonnet gravity.
It has been shown in the literature Jonas:2020pos that the no-boundary proposal remains robust even with the inclusion of higher-order terms in the Riemann curvature. These higher-order terms are expected to arise due to quantum gravity corrections. Our investigation into the Einstein-Gauss-Bonnet theory incorporates such higher-order terms, and the subsequent results align with this notion.
Acknowledgements
We are thankful to the organizers of the international conference QG@RRI 2023 for hosting the wonderful event and giving us a chance to present our work in it. We appreciate the warm hospitality and the vibrant platform the conference provided for exciting discussions. We also thank the anonymous referees for their helpful suggestions.
Appendix A Computation of Fluctuation Prefactor
In this appendix, we detail the computation of the fluctuation prefactor for all the boundary conditions discussed in the main text, following Ailiga:2023wzl ; doi:10.1142/7305 . Consider a one-dimensional quantum mechanical system described by the action
$S[q(t)]=\int_{0}^{1}\,\frac{m\dot{q}^{2}}{2}-V(q)\,dt.$ (115)
The path integral we wish to compute is when the Dirichlet boundary condition is imposed on both boundaries (DBC–DBC). The path integral is
$\bar{G}[q_{f};q_{i}]_{DD}=\int_{q(0)=q_{i}}^{q(1)=q_{f}}\mathcal{D}q(t)e^{iS[q(t)]}.$ (116)
The above path integral can be computed exactly for a quadratic potential of the form $V(q)=Aq^{2}+Bq+C$, where, $A,B$, and $C$ are $q$-independent constants. For this potential, the exact path integral acquires the form doi:10.1142/7305
$\bar{G}[q_{f};q_{i}]_{DD}=\mathcal{F}_{DD}e^{iS[q_{cl}(t)]}$ (117)
where,
$\mathcal{F}_{DD}=\left[-\frac{1}{2\pi i\hbar}\frac{\partial^{2}S[q_{cl}(t)]}{\partial q_{i}\partial q_{f}}\right]^{1/2}$ (118)
is the Van Vleck-Pauli-Morette fluctuation determinant for the DBC-DBC case, and $S[q_{cl}(t)]$, is the on-shell action corresponding to the classical solution $q_{cl}(t)$, that satisfies the on-shell equation
$m\ddot{q}_{cl}=-V^{\prime\prime}(q_{cl})$ (119)
subject to the Dirichlet-Dirichlet boundary condition.
Utilizing the above result, we proceed to compute the path integral with Neumann boundary condition (NBC) on the initial hypersurface $p(0)=p_{i}$, and Dirichlet boundary condition on the final hypersurface, $q(1)=q_{f}$. The path integral we wish to evaluate is
$\bar{G}[p_{i};q_{f}]_{ND}=\int_{p(0)=p_{i}}^{q(1)=q_{f}}\mathcal{D}q(t)e^{iS[q(t)]}$ (120)
where, $p(t)$ is the canonical momentum associated to $q(t)$. The above integral again can be evaluated exactly using the following relation Ailiga:2023wzl ; doi:10.1142/7305 ,
[…]
Extending the above definitions and relations to the minisuperspace quantum cosmology, eq. (18), and using the relations in eq. (47), the expression for the DBC–DBC path integral reads Feldbrugge:2017kzv
$\begin{split}\bar{G}_{\rm DBC}[q_{f},t=1;q_{i},t=0]=\sqrt{\frac{3i}{4\pi\hbar N_{c}}}\exp\left[\frac{i}{\hbar}\left\{\frac{N_{c}^{3}\Lambda^{2}}{36}+N_{c}\left(-\frac{\Lambda}{2}(q_{i}+q_{f})+3k\right)\right.\right.\\ \left.\left.+\frac{1}{N_{c}}\left(-\frac{3}{4}(q_{f}-q_{i})^{2}\right)\right\}\right],\end{split}$ (122)
where the exponent corresponds to the on-shell action for the DBC-DBC case, and the prefactor is the Van Vleck-Pauli-Morette fluctuation determinant ($\mathcal{F}_{DD}$). Note that the prefactor matches exactly with the expressions obtained in Ailiga:2023wzl . Furthermore, applying the relation, eq. (121, the path integral for Neumann - Dirichlet boundary conditions is given by
$\bar{G}_{\rm NBC}[q_{f},t=1;\pi_{i},t=0]=\exp\biggl{[}\frac{i}{\hbar}\biggl{\{}\frac{\Lambda^{2}N_{c}^{3}}{9}-\frac{\Lambda\pi_{i}N_{c}^{2}}{3}+\left(3k-\Lambda q_{f}+\frac{\pi_{i}^{2}}{3}\right)N_{c}+q_{f}\pi_{i}\biggr{\}}\biggr{]}\,.$ (123)
Here, as in the DBC–DBC case, the prefactor and exponent correspond to the fluctuation determinant and the on-shell action, respectively. It is crucial to note that the prefactor multiplied to the exponent is unity,
$\mathcal{F}_{ND}=1,$ (124)
which is the fluctuation determinant for the Neumann-Dirichlet case. This agrees with the eq. (49) in the main text and also align with the results obtained in Ailiga:2023wzl ; doi:10.1142/7305 .
One may extend the above analysis for the most general linear Robin boundary condition (RBC) on the initial hypersurface and the Dirichlet boundary condition (DBC) on the final hypersurface. The path integral we wish to evaluate is
$\bar{G}[q_{f};P_{i},\beta]_{RD}=\int_{p(0)+\beta q(0)=P_{i}}^{q(1)=q_{f}}\mathcal{D}q(t)e^{iS[q(t)]}.$ (125)
Again, the above path integral can be done exactly using the following relation Ailiga:2023wzl ,
[…]
Therefore, utilizing the result in eq. (122), the exact expression for the RBC-DBC path integral reads
$\begin{split}\bar{G}_{\rm RBC}[q_{f},t=1;P_{i},t=0]=\frac{1}{\sqrt{1+\frac{2}{3}N_{c}\beta}}\exp\left[\frac{1}{18(3+2N_{c}\beta)}\biggl{\{}\beta\Lambda^{2}N_{c}^{4}+6\Lambda^{2}N_{c}^{3}\right.\\ +\left.N_{c}^{2}\{108\beta k-18\Lambda(P_{i}+\beta q_{f})\}+18N_{c}\left\{9k+P_{i}^{2}+q_{f}\left(\beta^{2}q_{f}-3\Lambda\right)\right\}+54P_{i}q_{f}\biggr{\}}\right].\end{split}$ (127)
where the quantity multiplied to the exponent is the fluctuation prefactor,
$\mathcal{F}_{RD}=\frac{1}{\sqrt{1+\frac{2}{3}N_{c}\beta}}.$ (128)
In the $\beta\rightarrow 0$ limit (NBC-DBC), it goes to eq. 124.